### v1

# A Stochastic Gradient Descent Method with Global Convergence for Minimizing Nearly Convex Functions

Abstract.
This paper proposes a stochastic gradient descent method with an adaptive Gaussian noise term for minimizing nonconvex differentiable functions. The noise term in the algorithm, independent of the gradient, is determined by the difference between the function value at the current step and a lower bound estimate of the optimal value. In both probability space and state space, our theoretical analysis shows that for a class of nonconvex functions, represented by nearly convex functions, the proposed algorithm converges linearly to a certain neighborhood of the global optimal solution whose diameter depends on the variance of the gradient and the deviation between the estimated lower bound and the optimal value. Specifically, when full gradient information is utilized and the sharp lower bound of the objective function is available, the algorithm converges linearly to the global optimal solution. Furthermore, we propose a double-loop method that alternatively updates the lower bound estimate of the optimal value and the sequence, achieving the convergence to a neighborhood of the global optimal solution depending only on the variance of the gradient, provided that the lower bound estimate is asymptotically accurate. Numerical experiments on several concrete problems demonstrate the effectiveness of the proposed algorithm and validate the theoretical findings.
Key words and phrases:
stochastic gradient descent, global optimization, linear convergence, nearly convex functions
2020 Mathematics Subject Classification:
Primary 65K05, 90C26; Secondary 90C06
The research was supported partly by the National Key R&D Program of China (No. 2021YFA001300), the National Natural Science Foundation of China (Nos. 12271291, 12271150), the Hunan Provincial Natural Science Foundation of China (No. 2023JJ10001), and The Science and Technology Innovation Program of Hunan Province (No. 2022RC1190).
1. Introduction
Gradient descent (GD) is a classical method for solving the unconstrained nonlinear optimization problem
(1.1) $\min_{x\in^{d}}f(x),$
where d is the $d$-dimensional Euclidean space and the objective function $f$ is differentiable but not necessarily convex. Given an initial point $x_{0}\in^{d}$, it progressively calculates
$x_{t+1}:=x_{t}-\alpha_{t}\nabla f(x_{t}),\quad t=0,1,\ldots,$
until some stopping criterion is satisfied, where $\alpha_{t}>0$ is the step size (or the learning rate in the machine learning community).
The convergence properties of GD have been extensively studied when additional properties are imposed on $f$, such as the strong convexity and the Lipschitz continuity of the gradient. However, when $f$ is not convex, the iterates $\{x_{t}\}$ may easily become trapped around at local optimal solutions or stationary points, especially when the initial point $x_{0}$ is not intentionally chosen. One popular approach to escaping from non-optimal stationary points is to add a noise term that produces random perturbations to the GD iterations, i.e., setting
(1.2) $x_{t+1}:=x_{t}-\alpha_{t}\nabla f(x_{t})+\sigma_{t}\xi_{t},\quad t=0,1,\ldots,$
where $\xi_{t}$ is a random variables following a certain distribution, and $\sigma_{t}$ is the strength of the noise. A commonly used choice is to set $\xi_{t}\sim\mathcal{N}\left(0,\frac{1}{d}I_{d}\right)$, that is, the standard $d$ multivariate Gaussian distribution with 00 and $\frac{1}{d}I_{d}$ being the mean and the covariance matrix, respectively. Here, $I_{d}$ is the $d\times d$ identity matrix. In this setting, finding an appropriate decay rate of the noise strength $\sigma_{t}$ is important for the convergence to the global minimum, as discussed in [3, 8, 9, 10, 13]. When the step size $\alpha_{t}$ tends to zero, (1.2) approaches a continuous limit described by the stochastic differential equation
(1.3) $\mathrm{d}x=-\nabla f(x)\mathrm{d}t+\sigma(x,t)\mathrm{d}w,$
where $x:\to^{d}$ is an unknown differentiable function with respect to $t$, $w$ is a standard Brownian motion, and $\sigma$ is a real-valued function. The process $x$ in (1.3) is called the diffusion process, which can be used to study the behavior of the stochastic gradient descent (SGD) method. In general, it has been shown in [3, 9] that the critical decay rate of the noise strength $\sigma(x,t)$ is $\mathcal{O}(1/\sqrt{\ln t})$. Moreover, denote $G$ as the domain of attraction for a local minimum $\hat{x}$ [4], fix $\sigma(x,t)\equiv\sigma$ in (1.3), and let $x(0)\in G$ be the starting point. Then, according to the large deviation principle [4], the stopping time $\tau_{\sigma}$ that the diffusion process reaches the boundary $\partial G$ of $G$ satisfies
$\tau_{\sigma}\sim\exp\left(\frac{2}{\sigma^{2}}\inf_{x\in\partial G}(f(x)-f(\hat{x}))\right)\ \text{ as }\ \sigma\rightarrow 0^{+},$
which indicates that the time it takes for the process $x$ to escape from a local minimum could be exponentially long.
In a recent work [5], the authors proposed an adaptive scheme to adjust $\sigma_{t}$ in (1.2) that depends on $f(x_{t})$, i.e., the objective value in the current iteration. Specifically, given a monotonically decreasing sequence $\{f_{t}\}$, if $x_{t}$ lies in the sublevel set of $f$ determined by $f_{t}$, it accepts $x_{t+1}$ calculated by (1.2). Otherwise, it randomly samples $x_{t+1}$ from the standard multivariate Gaussian distribution. Under some assumptions about $f$, the algebraic convergence in both probability space and state space has been established in [5], which substantially improves the logarithmic convergence rate in [3, 9]. More recently, a derivative-free global optimization method was proposed in [6] with a similar algebraic convergence rate corresponding to $\alpha_{t}\equiv 0$ but with a carefully chosen scheme of $\sigma_{t}$ in (1.2).
In this work, we propose a new perturbed GD scheme in the framework of (1.2). Specifically, with $f_{\textsc{lb}}$ being a lower bound of $f$, the noise strength $\sigma_{t}$ is chosen so that it depends on the square root of the difference between $f(x_{t})$ and $f_{\textsc{lb}}$. Denote $x^{*}$ to be the unique global minimum point of $f$ and $f^{*}=f(x^{*})$. We prove the linear convergence of iterations $\{x_{t}\}$ to $x^{*}$ by setting $f_{\textsc{lb}}=f^{*}$ under the assumption that $f$ is a $(\alpha,L)$-nearly convex function, that is $L$-smooth with respect to $x^{*}$ and can be approximated by an $\alpha$-one point strongly convex function (cf. Section 2.1 for the details). Moreover, the proposed perturbation strategy also applies to the SGD method.
SGD and its variants are widely used optimization algorithms for training deep neural networks, especially when the full gradient is not accessible or too expensive to compute, e.g. when minimizing an expected risk given by $f(x):=\mathbb{E}_{z}\left[{}F(x,z)\right]$, where $z$ is a random vector drawn from an unknown distribution. SGD updates the sequence $\{x_{t}\}$ iteratively using the rule:
(1.4) $x_{t+1}:=x_{t}-\eta_{t}{\rm SG}(x_{t}),\quad t=0,1,\ldots.$
Here, $\eta_{t}$ is the learning rate and ${\rm SG}(x)$ is the stochastic gradient that satisfies $\mathbb{E}\left[{}{\rm SG}(x)\right]=\nabla f(x)$. Note that when using GD to minimize the expected risk $\mathbb{E}_{z}\left[{}F(x,z)\right]$ one may consider using the empirical risk $\frac{1}{n}\sum_{i=1}^{n}F(x,z_{i})$ as a surrogate. When $F(x,z)$ is strongly convex with respect to $x$ and $n$ is sufficiently large, it has been proven in [17, 2] that SGD with ${\rm SG}(x_{t})=\nabla_{x}F(x,z_{i})$ with $i$ being randomly chosen reaches an $\epsilon$-excess error (the error between the expected risk $f(x)$ and the optimal value $f^{*}$) with fewer gradient calls than GD in an asymptotic sense. When $f$ is not convex, the noise induced by ${\rm SG}(x)$ is critical for the provable ability of SGD to escape from saddle points [7, 11] or sharp minima [12, 14]. Like the parameter $\alpha_{t}$ in (1.2), the learning rate $\eta_{t}$ controls the noise strength and should be carefully tuned in practice [18]. However, to the best of our knowledge, the quantitative analysis for the convergence to the global minimum point is still under exploration. Combing (1.2) with (1.4), we conduct the convergence analysis for the update scheme:
(1.5) $x_{t+1}:=x_{t}-\eta_{t}{\rm SG}(x_{t})+\sigma_{t}\xi_{t},\quad t=0,1,\ldots,$
where $\sigma_{t}:=\sqrt{\eta_{t}s(f(x_{t}-\eta_{t}{\rm SG}(x_{t}))-f_{\textsc{lb}})}$ and $s>0$ is a hyper-parameter (cf. Section 3 for the details). If the objective function $f$ is $(\alpha,L)-$nearly convex, we show that the update scheme (1.5) will reach an $\epsilon$-neighborhood of global minimum point linearly in both probability space and state space, where the radius $\epsilon$ depends on the variance of the stochastic gradient ${\rm SG}(x)$ and the difference between $f^{*}$ and $f_{\textsc{lb}}$.
Additionally, we introduce a double-loop version of (1.5) (including (1.2)) which adaptively estimates $f_{\textsc{lb}}$ such that the gap between $f^{*}$ and $f_{\textsc{lb}}$ decreases below $\epsilon$ with high probability. In particular, for the double-loop version of (1.2) the iterates can find a point in the neighborhood with radius $\epsilon$ of the global minimum point in $\mathcal{O}(\ln^{2}(1/\epsilon))$ steps with high probability. Our results demonstrate that a specific class of nonconvex functions, called nearly convex functions, can be efficiently and effectively solved using the function value-dependent noise term in the proposed SGD method. It is worth mentioning that a nearly convex function can be nonconvex and contains infinitely many saddle points and local minima, as illustrated in Example 2.1.
The rest of this work is organized as follows. In Section 2, we provide the notation and summarize some preliminary results. In Section 3 we present the main algorithms and their convergence results. The proofs of the results in this paper are given in Section 4 and the numerical experiments are presented in Section 5. Finally, we conclude the work in Section 6.
2. Notation and preliminaries
[…]
For a continuously differentiable function $f:^{d}\to$, we use $f^{\prime}(x)$ and $\nabla f(x)$ to denote its Jacobian and gradient, respectively.
2.1. Nearly convex functions
In this part, we define and discuss a nonconvex function class called nearly convex functions. We first propose the definition of the one-point strongly convex functions and the nearly convex functions.
Definition 2.1.
The set of one-point strongly convex functions with $\alpha>0$ at $\bar{x}\in^{d}$ for $s\in$ is defined by
$\mathcal{P}_{\alpha}(\bar{x},s)=\{\bar{f}\in\mathcal{C}^{1}(^{d})\mid\bar{f}(\bar{x})=s,\langle\nabla\bar{f}(x),x-\bar{x}\rangle\geq\alpha\left\|x-\bar{x}\right\|^{2},\forall x\in^{d}\}.$
Note that for any $\bar{f}\in\mathcal{P}_{\alpha}(\bar{x},s)$, $\bar{x}$ is the unique global minimum point of $\bar{f}$ with the optimal value $s$ (see (a) in Lemma 2.1). Moreover, given a continuous differentiable function $f$, we define $\beta(f,\bar{x},\alpha)$ that measures the distance from $f$ to $\mathcal{P}_{\alpha}(\bar{x},f(\bar{x}))$.
Definition 2.2.
Given $\alpha>0$, $\bar{x}\in^{d}$, and $f\in\mathcal{C}^{1}(^{d})$, define the constant
$\beta(f,\bar{x},\alpha):=\inf_{\bar{f}\in\mathcal{P}_{\alpha}(\bar{x},f(\bar{x}))}\sup_{x\neq\bar{x}}\frac{2\left\lvert f(x)-\bar{f}(x)\right\rvert}{\left\|x-\bar{x}\right\|^{2}}.$
Remark 1.
Note that $\beta(f,\bar{x},\alpha)\leq C$ is equivalent to the following statement: $\forall n\in\mathbb{N}^{+}$, there exists $\bar{f}_{n}\in\mathcal{P}_{\alpha}(\bar{x},f(\bar{x}))$ such that
$\left\lvert f(x)-\bar{f}_{n}(x)\right\rvert\leq\left(\frac{C}{2}+\frac{1}{n}\right)\left\|x-\bar{x}\right\|^{2},\quad\forall x\in^{d}.$
Definition 2.3.
Let $L\geq\alpha>0$. We say a continuously differentiable function $f:^{d}\to$ is $(\alpha,L)$-nearly convex if the following three conditions hold simultaneously:
  1. (a)
    $f$ admits a global minimizer $x^{*}\in^{d}$;
  2. (b)
    the function $f$ satisfies
    (2.1) $\left\|\nabla f(x)\right\|\leq L\left\|x-x^{*}\right\|,\quad\forall x\in^{d};$
  3. (c)
    the distance $\beta(f,x^{*},\alpha)$ satisfies
    (2.2) $\beta(f,x^{*},\alpha)\leq\frac{1}{4}\sqrt{\frac{\alpha^{5}}{dL^{3}}}.$
We provide the following remark that explains the motivation behind this work, focusing on the study of nearly convex functions, and presenting their essential properties.
Remark 2.
Assume that the stochastic gradient oracle takes the form of ${\rm SG}(x)=\nabla f(x)+\omega$, where $\omega$ is a zero-mean random noise term that may depend on $x$, SGD (1.4) can be rewritten as
$x_{t+1}:=x_{t}-\eta_{t}\nabla f(x_{t})-\eta_{t}\omega_{t},\quad t=0,1,\ldots.$
Here, $\omega_{1},\omega_{2},\cdots$ are independent. By following from [12], we let $y_{t}:=x_{t}-\eta_{t}\nabla f(x_{t})$, so that $x_{t+1}=y_{t}-\eta_{t}\omega_{t}$. Then the above SGD iterations can be interpreted as updating $\{y_{t}\}$ via
(2.3) $y_{t+1}:=y_{t}-\eta_{t}\omega_{t}-\eta_{t}\nabla f(y_{t}-\eta_{t}\omega_{t}).$
Taking expectation with respect to $\omega_{t}$ in (2.3), one obtains
(2.4) $\mathbb{E}_{\omega_{t}}\left[{}y_{t+1}\right]=y_{t}-\eta_{t}\nabla\mathbb{E}_{\omega_{t}}\left[{}f(y_{t}-\eta_{t}\omega_{t})\right],$
where $\mathbb{E}_{\omega_{t}}\left[{}f(y_{t}-\eta_{t}\omega_{t})\right]:=\int f(y_{t}-\eta_{t}\omega)\mathrm{d}\mathbb{P}(\omega)$ can be viewed as the convolution of the function $f$ and the noise $\omega_{t}$. By defining $\hat{f}(y):=\mathbb{E}_{\omega_{t}}\left[{}f(y-\eta_{t}\omega_{t})\right]$, one has from (2.4) that $\mathbb{E}_{\omega_{t}}\left[{}y_{t+1}\right]$ is obtained by performing a single GD step on $\hat{f}(y)$ from $y_{t}$. Moreover, since the convolution has the smoothing effect, it may smooth the sharp local minima and intuitively explains how SGD helps us escape from the local minima.
Thus, if the function $f(x):=\bar{f}(x)+\epsilon(x)$, where $\bar{f}(x)$ is one-point strongly convex and $\epsilon(x)$ is a small perturbation function, we conjecture that the convolution mentioned above will reduce the role of $\epsilon(x)$ and the SGD step is mainly inducted by $\bar{f}(x)$. In fact, this manuscript validates this idea if the perturbation is not too large, as shown in Definition 2.3.
Lemma 2.1.
Let $f:^{d}\to$ be $(\alpha,L)$-nearly convex and $x^{*}$ be a global minimizer of $f$ with $f^{*}=f(x^{*})$. Then the following statements are valid.
  1. (a)
    For any function $\bar{f}\in\mathcal{P}_{\alpha}(x^{*},f^{*})$ , one has
    $\bar{f}(x)-f^{*}\geq\frac{\alpha}{2}\left\|x-x^{*}\right\|^{2}.$
  2. (b)
    One has $\beta:=\beta(f,x^{*},\alpha)<\alpha$ and
    (2.5) $\frac{\alpha-\beta}{2}\left\|x-x^{*}\right\|^{2}\leq f(x)-f(x^{*})\leq\frac{L}{2}\left\|x-x^{*}\right\|^{2}.$
    Moreover, $x^{*}$ is the unique global minimizer of $f$.
Proof.
  1. (a)
    Fix any $x\in^{d}$, and define the uni-variate function $g(t):=\bar{f}(x^{*}+t(x-x^{*}))$, so that $g(0)=\bar{f}(x^{*})$ and $g(1)=\bar{f}(x)$. Since $\bar{f}\in\mathcal{P}_{\alpha}(x^{*},f^{*})$, by definition 2.1 it satisfies $\langle\nabla\bar{f}(x),x-x^{*}\rangle\geq\alpha\left\|x-x^{*}\right\|^{2}$ for all $x\in\mathds{R}^{d}$. Then
    $g^{\prime}(t)=\langle\nabla\bar{f}(x^{*}+t(x-x^{*})),x-x^{*}\rangle\geq\alpha t\left\|x-x^{*}\right\|^{2}\quad\forall x\in^{d}.$
    Therefore, it comes from $\bar{f}^{*}=f^{*}$ that
    $\bar{f}(x)=g(1)=g(0)+\int_{0}^{1}g^{\prime}(t)\mathrm{d}t\geq f^{*}+\frac{\alpha}{2}\left\|x-x^{*}\right\|^{2}\quad\forall x\in^{d}.$
  2. (b)
    From definition 2.2, we know that for any $n\in\mathbb{N}^{+}$, there exists $\bar{f}_{n}\in\mathcal{P}_{\alpha}(x^{*},f(x^{*}))$ satisfying that $\left\lvert f(x)-\bar{f}_{n}(x)\right\rvert\leq\left(\frac{\beta}{2}+\frac{1}{n}\right)\left\|x-x^{*}\right\|^{2}$ for all $x\in^{d}$. From (a) we have
    $\bar{f}_{n}(x)-f(x^{*})\geq\frac{\alpha}{2}\left\|x-x^{*}\right\|^{2}\quad\forall x\in^{d}.$
    Thus, it holds that
    $f(x)-f(x^{*})=f(x)-\bar{f}_{n}(x)+\bar{f}_{n}(x)-f(x^{*})\geq\left(\frac{\alpha-\beta}{2}-\frac{1}{n}\right)\left\|x-x^{*}\right\|^{2}\quad\forall x\in^{d}.$
    We obtain the inequality on the left side using the limit with $n\rightarrow+\infty$. Similarly, according to (2.1), one has for any $x\in^{d}$,
    $f(x)-f(x^{*})$ $=\int_{0}^{1}\langle\nabla f(x^{*}+t(x-x^{*})),x-x^{*}\rangle\mathrm{d}t\leq\int_{0}^{1}Lt\left\|x-x^{*}\right\|^{2}\mathrm{d}t=\frac{L}{2}\left\|x-x^{*}\right\|^{2}.$
    Finally, since $\beta<\alpha$ by (2.2), one has $f(x)>f(x^{*})$ for any $x\neq x^{*}$.
Next, we give two examples of nearly convex functions that are nonconvex.
Example 2.1.
For $n,k\in\mathbb{N}^{+}$, define the functions
$f_{n,k}(x):=\frac{1}{2}x^{2}-\left(1+\frac{1}{k}\right)\int_{0}^{x}t(\sin t)^{2n}\mathrm{d}t,\quad x\in.$
Note that each $f_{n,k}$ is continuously differentiable with $f^{\prime}_{n,k}(x)=x(1-(1+k^{-1})(\sin x)^{2n})$. Moreover, the stationary points of $f_{n,k}$ are given by $x_{j}^{\pm}=j\pi\pm\arcsin{\left(\frac{k}{k+1}\right)^{\frac{1}{2n}}},j\in\mathbb{N}$, together with the origin.
Now, we show that for any $k\in\mathbb{N}^{+}$, there exists $n_{k}\in\mathbb{N}^{+}$, such that $f_{n,k}(x)$ is $(\frac{21}{25},1)$-nearly convex whenever $n\geq n_{k}$. Define $\bar{f}(x):=\frac{21}{50}x^{2}$, one has that $\bar{f}(x)$ is $\frac{21}{25}$-one point strongly convex. Moreover, it is easy to see that $\left\lvert f_{n,k}^{\prime}(x)\right\rvert\leq\left\lvert x-x^{*}\right\rvert$, i.e., $f_{n,k}^{\prime}(x)$ is Lipschitz continuous with $L=1$ being the corresponding modulus. Define $g_{n}(x):=\int_{0}^{x}(\sin t)^{2n}\mathrm{d}t$. For $x\in[0,\frac{\pi}{2}]$, one has $g_{n}^{\prime\prime}(x)=2n(\sin x)^{2n-1}\cos x\geq 0$, which implies that $g$ is convex over $\left[0,\frac{\pi}{2}\right]$. Thus, one can get
$\begin{array}[]{ll}g_{n}(x)\leq\frac{2}{\pi}g_{n}\left(\frac{\pi}{2}\right)x\leq\frac{2}{\pi}g_{n}(\pi)x,\quad\forall x\in\left[0,\frac{\pi}{2}\right],\end{array}$
For $x\in[\frac{\pi}{2},\pi]$, one has $g_{n}(x)\leq g_{n}(\pi)\leq\frac{2}{\pi}g_{n}(\pi)x$. Then, for any $x\geq 0$, by writing $x=y+j\pi$ with $y\in[0,\pi]$ and $j\in\mathbb{N}^{+}$, one can get
$\begin{array}[]{l}g_{n}(x)=g_{n}(y+j\pi)=jg_{n}(\pi)+g_{n}(y)\leq\frac{2}{\pi}g_{n}(\pi)(j\pi+y)=\frac{2}{\pi}g_{n}(\pi)x.\end{array}$
Consequently, one has $\left\lvert g_{n}(x)\right\rvert\leq\frac{2}{\pi}g_{n}(\pi)\left\lvert x\right\rvert$ for all $x\in$. Note that for any $n\in\mathbb{N}^{+}$ one has
$g_{n}(\pi)=\int_{0}^{\pi}(\sin t)^{2n}\mathrm{d}t=\frac{(2n-1)!!}{(2n)!!}\pi\quad\mbox{with}\quad\frac{(2n-1)!!}{(2n)!!}:=\frac{2n-1}{2n}\times\frac{2n-3}{2n-2}\times\cdots\times\frac{1}{2}.$
We have $g_{n}(\pi)$ decreasing along with $n$ and converges to zero. Then for any $k\in\mathbb{N}^{+}$, one can define $n_{k}$ by
(2.6) $n_{k}:=\min_{n}\left\{n\left|\frac{(2n-1)!!}{(2n)!!}\pi\leq\frac{2k\pi}{25(k+1)}\right.\right\}.$
Moreover, one has $g_{n}(\pi)\leq\frac{2k\pi}{25(k+1)},\;\forall n\geq n_{k}$. When $n\geq n_{k}$, it holds that
$0\leq\int_{0}^{x}t(\sin t)^{2n}\mathrm{d}t=\int_{0}^{\left\lvert x\right\rvert}\left\lvert t\right\rvert(\sin t)^{2n}\mathrm{d}t\leq\left\lvert x\right\rvert\left\lvert g_{n}(x)\right\rvert<\frac{2}{\pi}g_{n}(\pi)x^{2}\leq\frac{4k}{25(k+1)}x^{2}.$
Hence one can get
$-\frac{4}{25}x^{2}\leq f_{n,k}(x)-\frac{1}{2}x^{2}\leq 0,$
and, as a consequence,
$\left\lvert f_{n,k}(x)-\bar{f}(x)\right\rvert\leq\frac{2}{25}x^{2}<\frac{\alpha}{8}\sqrt{\frac{\alpha^{3}}{dL^{3}}}\left\|x-x^{*}\right\|^{2}.$
Recall $\alpha=\frac{21}{25}$ and $L=d=1$. This validates that $f_{n,k}$ is $(\frac{21}{25},1)$-nearly convex whenever $n\geq n_{k}$. Moreover, from (2.6), one can explicitly calculate that $n_{1}=199,n_{2}=112,n_{3}=89$, and $n_{4}=78$.
Example 2.2.
For $\varepsilon\in(0,1)$ and $R>0$, consider the function
(2.7) $f(x)=\begin{cases}\dfrac{1+\varepsilon\sin\left(2R\log\left\lvert x\right\rvert\right)}{2}x^{2},&x\neq 0,\\ 0,&x=0.\end{cases}$
Obviously, $f(x)$ is continuously differentiable at $x=0$ with $f^{\prime}(0)=0$. And we observe that $\left\lvert f(x)-\frac{1}{2}x^{2}\right\rvert\leq\frac{\varepsilon}{2}x^{2}$, this shows that $x^{*}=0$ is the unique global minimizer of $f$, and $\beta(f,x^{*},\alpha)\leq\varepsilon$ with $\alpha=1$. Next, we prove the following nearly convex properties of $f$:
  • When $\varepsilon\sqrt{1+R^{2}}<1$, $f$ is $(1-\varepsilon\sqrt{1+R^{2}},1+\varepsilon\sqrt{1+R^{2}})$-nearly convex.
  • When $\varepsilon\sqrt{1+R^{2}}\geq 1$ and $4\varepsilon\left(1+\varepsilon\sqrt{1+R^{2}}\right)^{\frac{3}{2}}\leq 1$, $f$ is $(1,1+\varepsilon\sqrt{1+R^{2}})$-nearly convex.
For the derivation of $f$, when $x\neq 0$, we obtain
$f^{\prime}(x)$ $=(1+\varepsilon\sin(2R\log\left\lvert x\right\rvert)+\varepsilon R\cos(2R\log\left\lvert x\right\rvert))\cdot x$
$=\left(1+\varepsilon\sqrt{1+R^{2}}\sin(2R\log\left\lvert x\right\rvert+\xi)\right)x,$
where the angle $\xi$ satisfies
$\sin\xi=\frac{R}{\sqrt{1+R^{2}}},\quad\cos\xi=\frac{1}{\sqrt{1+R^{2}}}.$
This leads to
$\left\lvert f^{\prime}(x)\right\rvert\leq\left(1+\varepsilon\sqrt{1+R^{2}}\right)\left\lvert x\right\rvert,$
which shows that $f$ in (2.7) satisfies the condition (2.1) with $L=1+\varepsilon\sqrt{1+R^{2}}$. When $\varepsilon\sqrt{1+R^{2}}<1$, one can see that
$\left\langle{\nabla f(x),x}\right\rangle\geq\left(1-\varepsilon\sqrt{1+R^{2}}\right)x^{2},$
which shows that $f\in\mathcal{P}_{1-\varepsilon\sqrt{1+R^{2}}}(0,0)$. In this case, $\beta(f,x^{*},1-\varepsilon\sqrt{1+R^{2}})=0$ and we have that $f$ is $(1-\varepsilon\sqrt{1+R^{2}},1+\varepsilon\sqrt{1+R^{2}})$-nearly convex. When $\varepsilon\sqrt{1+R^{2}}\geq 1$, we take $\alpha=1$ and get $\beta(f,x^{*},\alpha)\leq\varepsilon$. Hence the condition (2.2) holds if
$\varepsilon\leq\dfrac{1}{4\left(1+\varepsilon\sqrt{1+R^{2}}\right)^{\frac{3}{2}}}.$
In this case, we know that $f$ is $(1,1+\varepsilon\sqrt{1+R^{2}})$-nearly convex.
Remark 3.
Note that for most other classes of nonconvex functions, such as those satisfying the Polyak-Lojasiewicz (PL) condition (cf. [1, 15]), it is typically required $\varepsilon\sqrt{1+R^{2}}<1$ to establish convergence proofs. However, in the framework of nearly convex functions, we can handle the more challenging cases where $\varepsilon\sqrt{1+R^{2}}\geq 1$.
3. Gaussian noise descent algorithms
In this part, we propose and analyze a Gaussian noise descent (GND) algorithm, as well as a double-loop Gaussian noise descent (DL-GND) algorithm, for solving (1.1). We make the following assumption on the stochastic gradient oracle.
Input: The initial point $x_{0}\in^{d}$, the stochastic gradient oracle ${\rm SG}(x)$, the learning rate $\eta>0$, the noise factor $s>0$, a lower bound $f_{\textsc{lb}}$ of $f$, and the total number of iterations $T$.
Output: $x_{t_{*}}$.
1 for $t=0,\ldots,T-1$ do
2       $x_{t+\frac{1}{2}}\leftarrow x_{t}-\eta{\rm SG}(x_{t})$;
3
4      $\sigma_{t}\leftarrow\sqrt{\eta s\left(f\left(x_{t+\frac{1}{2}}\right)-f_{\textsc{lb}}\right)}$;
5
6      $x_{t+1}\leftarrow x_{t+\frac{1}{2}}-\sigma_{t}\xi_{t}$ with $\xi_{t}\in\mathcal{N}\left(0,\dfrac{1}{d}I_{d}\right)$;
7
$t_{*}\leftarrow\min\left\{\operatorname*{arg\,min}\limits_{0\leq t\leq T}f\left(x_{t}\right)\right\}$.
Algorithm 1 A Gaussian noise Descent (GND) algorithm for solving (1.1)
Assumption 3.1.
There exists a constant $r\geq 0$ such that the stochastic gradient oracle ${\rm SG}(x)$ satisfies
$\mathbb{E}\left[{}{\rm SG}(x)\right]=\nabla f(x)\quad\mbox{and}\quad\mathbb{E}\left[{}\left\|{\rm SG}(x)-\nabla f(x)\right\|^{2}\right]\leq r^{2}.$
Note that when the objective function $f:^{d}\to$ in (1.1) is $(\alpha,L)$-nearly convex and a learning rate $\eta<\frac{2\alpha}{L^{2}}$ is used in Algorithm 1, one can define the two constants
(3.1) $\lambda:=2\alpha-\eta L^{2}>0,$
(3.2) $b:=\frac{\eta r^{2}}{2\alpha-\eta L^{2}}+\frac{5\eta(2\alpha-\eta L^{2})+14}{42L}(f^{*}-f_{\textsc{lb}}),$
where $f^{*}:=f(x^{*})$ is the minimum value of $f$, $f_{\textsc{lb}}\leq f^{*}$ is a lower bound estimation of $f$, and $r>0$ is the constant such that 3.1 holds. Next, present the global convergence of Algorithm 1 for the case that $f$ is nearly convex and Assumption 3.1 holds. For ease of presenting the most essential conclusion of our results, we take $\eta=\frac{2\alpha}{5L^{2}},s=\frac{\lambda}{3L}$ in the following convergence theorem, and one can obtain similar results by using similar analysis (see remark 4 for more details).
Theorem 3.1.
Suppose that the objective function $f:^{d}\to$ in (1.1) is $(\alpha,L)$-nearly convex and that 3.1 holds. Define $\eta=\frac{2\alpha}{5L^{2}},s=\frac{\lambda}{3L}$, and let $f_{\textsc{lb}}\leq f^{*}$ be any lower bound estimation of $f$. Let $\lambda,b\geq 0$ be the constant defined by (3.1) and (3.2). Let $\{x_{t}\}$ be the sequence generated by Algorithm 1 with the specified $\eta,s,f_{\textsc{lb}}$. Define
$y_{t}:=x_{t}-\eta\nabla f(x_{t}),\quad t=0,\ldots,T-1.$
Then the following conclusions hold.
  1. (a)
    If $b=0$, for any $\varepsilon>0$ one has $\left\|y_{T}-x^{*}\right\|^{2}\leq\varepsilon$ with probability at least $1-\zeta$, provided that $\zeta\in(0,1)$ and
    $T\geq\frac{100}{\eta(2\alpha-\eta L^{2})}\ln\frac{\left\|y_{0}-x^{*}\right\|^{2}}{\zeta\varepsilon}.$
  2. (b)
    If $b>0$, there exists a certain $t\leq T$ satisfying $\left\|y_{t}-x^{*}\right\|^{2}\leq 200b$ with probability at least $1-\zeta$, provided that $\zeta\in(0,1)$ and
    $T\geq\frac{200}{\eta(2\alpha-\eta L^{2})}\ln\frac{\left\|y_{0}-x^{*}\right\|^{2}}{100b\zeta}.$
    Moreover, by letting $\beta:=\beta(f,x^{*},\alpha)<\alpha$, one has with probability at least $1-\zeta$ that
    (3.3) $f(x_{t_{*}})-f^{*}\leq\frac{100Lb}{(1-\eta L)^{2}},$
    (3.4) $\left\|x_{t_{*}}-x^{*}\right\|^{2}\leq\frac{200Lb}{(1-\eta L)^{2}(\alpha-\beta)},$
    (3.5) $\left\|y_{t_{*}}-x^{*}\right\|^{2}\leq\left(\frac{1+\eta L}{1-\eta L}\right)^{2}\frac{200Lb}{\alpha-\beta}.$
The proof of Theorem 3.1 are postponed to Section 4.2. Here we make the following remarks on Theorem 3.1 and its proof that will be given in Section 4.2.
Remark 4.
As will be observed from (4.14) in Section 4.2 (the proof of Theorem 3.1), the global linear convergence of Algorithm 1 still holds for $\eta<\frac{2\alpha}{L^{2}}$ , provided that a constant $s>0$ can be found such that $\beta=\beta(f,x^{*},\alpha)$ satisfies the following inequality:
(3.6) $\beta\sqrt{2d}\left(\sqrt{\frac{2}{\pi\eta s(\alpha-\beta)}}+1\right)\left(1+\eta sL\right)\leq 2\alpha-\eta L^{2}-\frac{sL}{2}.$
In our proof, we typically set $\eta=\frac{2\alpha}{5L^{2}}$ and $s=\frac{\lambda}{3L}$. In addition, (3.6) implies that $\eta$ cannot be arbitrarily small. However, when $f(x)$ is one-point strongly convex at $x^{*}$, meaning $\beta=0$, we can arbitrarily choose $\eta\in(0,\frac{2\alpha}{L^{2}})$ by letting $s=0$.
Remark 5.
If the full gradient is used and $f_{\textsc{lb}}=f^{*}$, 3.1 holds with $r=0$, which implies $b=0$. In this case, Theorem 3.1 shows that the sequence $\{y_{t}\}$ converges linearly to $x^{*}$. If $r>0$ and $f_{\textsc{lb}}=f^{*}$, Theorem 3.1 shows that $\{y_{t}\}$ converges linearly to a neighborhood of $x^{*}$, and the size of the neighborhood depends on the variance $r$ of the stochastic oracle. More specifically, if $f$ is one-point strongly convex at $x^{*}$, we can take $\eta=\epsilon$. Thus, with high probability, $\{y_{t}\}$ achieves a neighborhood $\mathcal{O}(\epsilon)$ of $x^{*}$ within $T=\mathcal{O}\left(\frac{1}{\epsilon}\ln\frac{1}{\epsilon}\right)$ iterations.
Note that, when implementing Algorithm 1, it is generally impossible to evaluate $f^{*}$ as a sharp lower bound of $f$, but estimating an alternative lower bound can be much easier. For example, many loss functions used in machine learning usually admit 00 as a trivial lower bound. Based on such an observation, we propose a double-loop version of Algorithm 1. For convenience, we abbreviate Algorithm 1 to
$x_{t_{*}}\leftarrow\operatorname{GND}({\rm SG},x_{0},\eta,s,f_{\textsc{lb}},T),$
where the inputs can be viewed as the variables of a function named as GND and the unique output can be viewed as the function value. The double-loop variant of Algorithm 1 is given as follows.
Input: The initial point $x_{0}\in^{d}$, the stochastic gradient oracle ${\rm SG}(x)$, the learning rate $\eta>0$, the noise factor $s>0$, a lower bound $f_{\textsc{lb}}^{0}$ of $f$, a combination coefficient $\gamma\in(0,1)$, a number $N$ of the outer loops, and the two numbers $T_{1}$ and $T_{2}$ for the total iterations in the first and second stages, respectively.
Output: $x_{\min}^{N}$.
1 $x_{\min}^{0}\leftarrow\operatorname{GND}({\rm SG},x_{0},\eta,s,f_{\textsc{lb}}^{0},T_{1})$ by implementing Algorithm 1;
2for $\nu=0$ to $\nu=N-1$ do
3       $f_{\textsc{lb}}^{\nu+1}\leftarrow(1-\gamma)f_{\textsc{lb}}^{\nu}+\gamma f\left(x_{\min}^{\nu}\right)$;
4
5      $x_{\min}^{\nu+1}\leftarrow\operatorname{GND}({\rm SG},x_{\min}^{\nu},\eta,s,f_{\textsc{lb}}^{\nu+1},T_{2})$ by implementing Algorithm 1.
Algorithm 2 A double-loop Gaussian noise descent (DL-GND) algorithm for solving (1.1)
In Algorithm 2, the outer loop mainly updates the lower bound estimate of the objective function, while the inner loop is similar to Algorithm 1. Moreover, we have the following result regarding the convergence properties of Algorithm 2.
Theorem 3.2.
Suppose that the objective function $f:^{d}\to$ in (1.1) is $(\alpha,L)$-nearly convex and 3.1 holds. Let $\eta=\frac{2\alpha}{5L^{2}},s=\frac{\lambda}{3L}$ and assume that $f_{\textsc{lb}}^{0}<f^{*}$. For any $\varepsilon>0$ and $\zeta\in(0,1)$, define
(3.7) $\begin{array}[]{ll}b^{0}:=\frac{\eta r^{2}}{\lambda}+\frac{5\eta\lambda+14}{42L}(f^{*}-f_{\textsc{lb}}^{0}),\quad b^{\varepsilon}:=\frac{\eta r^{2}}{\lambda}+\frac{5\eta\lambda+14}{42L}\varepsilon,\quad\mbox{and}\quad\gamma:=\frac{(1-\eta L)^{2}\varepsilon}{(1-\eta L)^{2}\varepsilon+100Lb^{\varepsilon}}.\end{array}$
where $\lambda$ is defined in (3.1). Denote $\beta=\beta(f,x^{*},\alpha)$. Suppose that the positive integers $N$, $T_{1}$, and $T_{2}$ in Algorithm 2 are chosen such that
(3.8) $\begin{array}[]{ll}N\geq\frac{1}{\gamma}\ln\left((1-\gamma)+\gamma\cdot\frac{f^{*}-f_{\textsc{lb}}^{0}}{\varepsilon}\right),\quad T_{1}\geq\frac{200}{\eta\lambda}\ln\frac{\left\|y_{0}-x^{*}\right\|^{2}}{100b^{0}\zeta^{\prime}},\quad\mbox{and}\quad T_{2}\geq\frac{200}{\eta\lambda}\ln\frac{2(1+\eta L)^{2}Lb^{0}}{(1-\eta L)^{2}(\alpha-\beta)b^{\varepsilon}\zeta^{\prime}}\end{array}$
with $\zeta^{\prime}:=\frac{\zeta}{N+1}$. Then, with probability at least $1-\zeta$, one has that the output $x^{N}_{\min}$ of Algorithm 2 satisfies
(3.9) $\left\|x_{\min}^{N}-x^{*}\right\|^{2}\leq\frac{200Lb^{\varepsilon}}{(1-\eta L)^{2}(\alpha-\beta)}.$
The proof of Theorem 3.2 will be given in Section 4.3. Here we make a remark on Theorem 3.2.
Remark 6.
In Algorithm 2, the total number of iterations to achieve the convergence bound in (3.9) has the estimations that
$T_{1}+NT_{2}=\begin{cases}\mathcal{O}\left(\ln^{2}\frac{1}{\varepsilon}\right)&\mbox{ if }r=0,\\ \mathcal{O}\left(\frac{1}{\varepsilon}\ln\frac{1}{\varepsilon}\right)&\mbox{ if }r>0.\end{cases}$
Therefore, if we use the full gradient $(r=0)$ in Algorithm 2, it will find a point in the neighborhood with radius $\epsilon$ of the global minima $x^{*}$ within $\mathcal{O}\left(\ln^{2}\frac{1}{\varepsilon}\right)$ steps with high probability.
4. Proofs of the main results
This section first presents preliminary results related to the probability theory and stochastic process and then provides the proofs for Theorems 3.1 and 3.2.
4.1. Preliminaries
Recall that the Gamma function on positive real numbers is defined by
$\Gamma(z):=\int_{0}^{\infty}t^{z-1}e^{t}\mathrm{d}t,\quad 0<z\in.$
The following Lemma is a basic result of probability theory.
Lemma 4.1.
Let $\xi\sim\mathcal{{}N}(0,\frac{1}{d}I_{d})$ be a random vector in d and $\Gamma(\cdot)$ be the Gamma function. One has that
(4.1) $\begin{array}[]{ll}\mathbb{E}\left[{}\left\|\xi\right\|\right]=\frac{\sqrt{2}}{\sqrt{d}}\frac{\Gamma\left(\frac{d+1}{2}\right)}{\Gamma\left(\frac{d}{2}\right)},\quad\mathbb{E}\left[{}\left\|\xi\right\|^{2}\right]=1,\quad\mathbb{E}\left[{}\left\|\xi\right\|^{3}\right]=\frac{\sqrt{2}}{\sqrt{d}}\frac{\Gamma\left(\frac{d+1}{2}\right)}{\Gamma\left(\frac{d}{2}\right)}\left(1+\frac{1}{d}\right)\ \,\mbox{and}\quad\mathbb{E}\left[{}\left\|\xi\right\|^{4}\right]=1+\frac{2}{d}.\end{array}$
Proof.
Note that $\|\sqrt{d}\xi\|$ is a random variable following a $\chi$-distribution with $d$ degrees of freedom. One can readily get (4.1) by examining the raw moments of the $\chi$-distribution. ∎
Based on the above Lemma, we have the following preliminary result, whose proof is given in Section A.1.
Lemma 4.2.
Let $\varepsilon:^{d}\to$ be a continuously differentiable function such that $\left\lvert\varepsilon(x)\right\rvert\leq\frac{l}{2}\left\|x-x^{*}\right\|^{2}$ with $l>0$ and $x^{*}\in^{d}$. Given any $x\in^{d}$ and $\delta>0$, one has for $\xi\sim\mathcal{{}N}(0,\frac{1}{d}I_{d})$ that
(4.2) $\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\varepsilon(x-\delta\xi),x-x^{*}\rangle\right]\right\rvert\leq l\sqrt{\frac{d}{2}}\frac{\left\|x-x^{*}\right\|}{\delta\sqrt{\pi}}\left(\left\|x-x^{*}\right\|^{2}+\delta^{2}\left(1+\frac{1}{d}\right)\right),$
(4.3) $\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\varepsilon(x-\delta\xi),\delta\xi\rangle\right]\right\rvert\leq l\sqrt{\frac{d}{2}}\left(\left\|x-x^{*}\right\|^{2}+\delta^{2}\left(1+\frac{1}{d}\right)\right),$
(4.4) $\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\varepsilon(x-\delta\xi),x-\delta\xi-x^{*}\rangle\right]\right\rvert\leq l\sqrt{\frac{d}{2}}\left(\frac{\left\|x-x^{*}\right\|}{\delta\sqrt{\pi}}+1\right)\left(\left\|x-x^{*}\right\|^{2}+\delta^{2}\left(1+\frac{1}{d}\right)\right).$
Note that Lemma 4.2 contains the specific results of the convolution effect discussed in Remark 2. The following lemma on stochastic processes with its proof given in Section A.2 will be used in the convergence analysis of the algorithm proposed in the next section.
Lemma 4.3.
[…]
Then, for any given nonnegative integer $M$ and real number $\ell>0$,
(4.5) $\mathbb{P}\left[{}\exists t\leq M,X_{t}<\ell\right]\geq 1-\left(\frac{b+\theta\ell}{b+\ell}\right)^{M}\frac{B}{\ell}\quad\mbox{with}\quad B:=\mathbb{E}\left[{}X_{0}\mathds{1}\{X_{0}\geq\ell\}\right]\geq 0.$
4.2. Proof of Theorem 3.1
The following Lemma plays a key role in the convergence analysis.
Lemma 4.4.
Suppose that the function $f:^{d}\to$ in (1.1) is $(\alpha,L)$-nearly convex with $x^{*}\in^{d}$ being one of its global minimizers. Define $f^{*}=f(x^{*})$ and let $f_{\textsc{lb}}\leq f^{*}$ be a lower bound estimation of $f$. Let $\lambda>0$ and $b\geq 0$ be the two constants defined by (3.2) with $\eta:=\frac{2\alpha}{5L^{2}}$. Define the constant $s:=\frac{\lambda}{3L}$ and the functions
$G(x):=x-\eta\nabla f(x)\quad\mbox{and}\quad\sigma(x):=\sqrt{\eta s\left(f(x)-f_{\textsc{lb}}\right)},\quad x\in^{d},$
Then, one has with $\xi\sim\mathcal{N}(0,\frac{1}{d}I_{d})$ that
(4.6) $\mathbb{E}_{\xi}\left[{}\left\|G(x-\sigma(x)\xi)-x^{*}\right\|^{2}\right]\leq\left(1-\frac{\eta\lambda}{100}\right)\left\|x-x^{*}\right\|^{2}+\frac{\eta\lambda(5\eta\lambda+14)}{42L}(f^{*}-f_{\textsc{lb}})\quad\forall x\in^{d}.$
Proof.
Note that
(4.7) $\left\|G(x-\sigma(x)\xi)-x^{*}\right\|^{2}=\left\|x-\sigma(x)\xi-\eta\nabla f(x-\sigma(x)\xi)-x^{*}\right\|^{2}$
$=\left\|x-\sigma(x)\xi-x^{*}\right\|^{2}+\left\|\eta\nabla f(x-\sigma(x)\xi)\right\|^{2}-2\eta\langle\nabla f(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle$
$\leq(1+\eta^{2}L^{2})\left\|x-\sigma(x)\xi-x^{*}\right\|^{2}-2\eta\langle\nabla f(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle,$
where the last inequality comes from (2.1) of Definition 2.3. By definition 2.2 we know that for any $n\in\mathbb{N}^{+}$, there exists $\bar{f}_{n}\in\mathcal{P}_{\alpha}(x^{*},f(x^{*}))$ satisfying that $\left\lvert f(x)-\bar{f}_{n}(x)\right\rvert\leq\left(\frac{\beta(f)}{2}+\frac{1}{n}\right)\left\|x-x^{*}\right\|^{2},\forall x\in^{d}$. We use $\beta=\beta(f)$ in the following. Denote $\epsilon_{n}=f-\bar{f}_{n}$, then we have $\left\lvert\epsilon_{n}(x)\right\rvert\leq\left(\frac{\beta}{2}+\frac{1}{n}\right)\left\|x-x^{*}\right\|^{2}$. From definition 2.1, $\langle\nabla\bar{f}_{n}(x),x-x^{*}\rangle\geq\alpha\left\|x-x^{*}\right\|^{2}$ for all $x\in\mathds{R}^{d}$. Therefore,
(4.8) $-2\eta\langle\nabla f(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle$
$=-2\eta\langle\nabla\bar{f}_{n}(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle-2\eta\langle\nabla\epsilon_{n}(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle$
$\leq-2\eta\alpha\left\|x-\sigma(x)\xi-x^{*}\right\|^{2}-2\eta\langle\nabla\epsilon_{n}(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle.$
Recall from (3.1) that $\lambda=2\alpha-\eta L^{2}>0$. Combining (4.7) and (4.8) obtains
$\left\|G(x-\sigma(x)\xi)-x^{*}\right\|^{2}\leq(1-\eta\lambda)\left\|x-\sigma(x)\xi-x^{*}\right\|^{2}-2\eta\langle\nabla\epsilon_{n}(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle.$
This inequality holds for any $n\in\mathbb{N}^{*}$. Therefore, by taking expectation with respect to $\xi$ and using (4.1) of Lemma 4.1, one gets
(4.9) $\mathbb{E}_{\xi}\left[{}\left\|G(x+\sigma(x)\xi)-x^{*}\right\|^{2}\right]$
$\leq(1-\eta\lambda)\mathbb{E}_{\xi}\left[{}\left\|x+\sigma(x)\xi-x^{*}\right\|^{2}\right]+2\eta\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\epsilon_{n}(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle\right]\right\rvert$
$=(1-\eta\lambda)\left(\left\|x-x^{*}\right\|^{2}+\sigma(x)^{2}\right)+2\eta\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\epsilon_{n}(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle\right]\right\rvert.$
Note we have $\left\lvert\epsilon_{n}(x)\right\rvert\leq\left(\frac{\beta}{2}+\frac{1}{n}\right)\left\|x-x^{*}\right\|^{2}$. Therefore, by taking $\delta=\sigma(x)$ in (4.4) of Lemma 4.2, one gets
(4.10) $\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\epsilon_{n}(x-\sigma(x)\xi),x-\sigma(x)\xi-x^{*}\rangle\right]\right\rvert$
$\leq\left(\beta+\frac{2}{n}\right)\sqrt{\frac{d}{2}}\left(\frac{\left\|x-x^{*}\right\|}{\sqrt{\pi}\sigma(x)}+1\right)\left(\left\|x-x^{*}\right\|^{2}+2\sigma(x)^{2}\right).$
Combine (4.9) and (4.10) gives
$\mathbb{E}_{\xi}\left[{}\left\|G(x+\sigma(x)\xi)-x^{*}\right\|^{2}\right]$
$\leq(1-\eta\lambda)\left(\left\|x-x^{*}\right\|^{2}+\sigma(x)^{2}\right)+2\eta\left(\beta+\frac{2}{n}\right)\sqrt{\frac{d}{2}}\left(\frac{\left\|x-x^{*}\right\|}{\sqrt{\pi}\sigma(x)}+1\right)\left(\left\|x-x^{*}\right\|^{2}+2\sigma(x)^{2}\right).$
Taking $n\rightarrow+\infty$ we get
(4.11) $\mathbb{E}_{\xi}\left[{}\left\|G(x+\sigma(x)\xi)-x^{*}\right\|^{2}\right]$
$\leq(1-\eta\lambda)\left(\left\|x-x^{*}\right\|^{2}+\sigma(x)^{2}\right)+\eta\beta\sqrt{2d}\left(\frac{\left\|x-x^{*}\right\|}{\sqrt{\pi}\sigma(x)}+1\right)\left(\left\|x-x^{*}\right\|^{2}+2\sigma(x)^{2}\right).$
Note that $f_{\textsc{lb}}\leq f^{*}$, so that by the definition of $\sigma(x)$ one has
$\sqrt{\eta s\left(f(x)-f^{*}\right)}\leq\sqrt{\eta s\left(f(x)-f_{\textsc{lb}}\right)}=\sigma(x)=\sqrt{\eta s\left(f(x)-f^{*}\right)+\eta s(f^{*}-f_{\textsc{lb}})}.$
According to (2.5) of Lemma 2.1, we conclude that
(4.12) $\sqrt{\frac{\eta s(\alpha-\beta)}{2}}\left\|x-x^{*}\right\|\leq\sigma(x)\leq\sqrt{\frac{\eta sL}{2}\left\|x-x^{*}\right\|^{2}+\eta s(f^{*}-f_{\textsc{lb}})}.$
Take (4.12) into (4.11), one has
(4.13) $\mathbb{E}_{\xi}\left[{}\left\|G(x+\sigma(x)\xi)-x^{*}\right\|^{2}\right]\leq(1-\eta\lambda)\left(\left(1+\frac{\eta sL}{2}\right)\left\|x-x^{*}\right\|^{2}+\eta s(f^{*}-f_{\textsc{lb}})\right)$
$\quad+\eta\beta\sqrt{2d}\left(\sqrt{\frac{2}{\pi\eta s(\alpha-\beta)}}+1\right)\left((1+\eta sL)\left\|x-x^{*}\right\|^{2}+2\eta s(f^{*}-f_{\textsc{lb}})\right).$
Denote $C_{1}:=\beta\sqrt{2d}\left(\sqrt{\frac{2}{\pi\eta s(\alpha-\beta)}}+1\right)$, and then we organize (4.13) as follows
(4.14) $\mathbb{E}_{\xi}\left[{}\left\|G(x+\sigma(x)\xi)-x^{*}\right\|^{2}\right]$
$\leq\left[1-\eta\left(\lambda\left(1+\frac{1}{2}\eta sL\right)-C_{1}\left(1+\eta sL\right)-\frac{1}{2}sL\right)\right]\left\|x-x^{*}\right\|^{2}$
$\quad+(1-\eta(\lambda-2C_{1}))\eta s(f^{*}-f_{\textsc{lb}}).$
By (2.2) we have $\beta=\beta(f)\leq\frac{1}{4}\sqrt{\frac{\alpha^{5}}{dL^{3}}}$, $L\geq\alpha$ and $d\geq 1$. One has $\beta\leq\frac{1}{4\sqrt{d}}\sqrt{\frac{\alpha^{5}}{L^{3}}}\leq\frac{\alpha}{4}$, so that $\frac{L}{\alpha-\beta}\leq\frac{4L}{3\alpha}$. Meanwhile, by $\eta=\frac{2\alpha}{5L^{2}}$ and $\lambda=2\alpha-\eta L^{2}=\frac{8\alpha}{5}$ one has $\eta\lambda=\frac{16\alpha^{2}}{25L^{2}}$. Note that
$\sqrt{\frac{2}{\pi\eta s(\alpha-\beta)}}=\sqrt{\frac{2}{\pi}\frac{3L}{\alpha-\beta}\frac{1}{\eta\lambda}}\leq 5\sqrt{\frac{L^{3}}{2\pi\alpha^{3}}},$
and it is easy to verify that $\frac{5}{\sqrt{\pi}}+\sqrt{2}<\frac{152}{35}$. Therefore,
(4.15) $C_{1}\leq\frac{\alpha}{4}\sqrt{\frac{\alpha^{3}}{L^{3}}}\left(5\sqrt{\frac{L^{3}}{\pi\alpha^{3}}}+\sqrt{2}\right)\leq\frac{\alpha}{4}\left(\frac{5}{\sqrt{\pi}}+\sqrt{2}\right)<\frac{38}{35}\alpha=\frac{19}{28}\lambda.$
Moreover, one has $\eta\lambda\leq\frac{16}{25}$, which, together with $s=\frac{\lambda}{3L}$, implies $1+\eta sL=\frac{\eta\lambda}{3}+1\leq\frac{91}{75}$. Consequently,
$C_{1}\left(1+\eta sL\right)<\frac{19}{28}\cdot\frac{91}{75}\lambda=\left(\frac{99}{100}-\frac{1}{6}\right)\lambda.$
Taking this and (4.15), together with $sL=\frac{\lambda}{3}$, we get
(4.16) $\lambda\left(1+\frac{1}{2}\eta sL\right)-C_{1}\left(1+\eta sL\right)-\frac{1}{2}sL\geq\lambda\left(1+\frac{\eta\lambda}{6}\right)-\left(\frac{99}{100}-\frac{1}{6}\right)\lambda-\frac{\lambda}{6}>\frac{\lambda}{100},$
(4.17) $(1-\eta(\lambda-2C_{1}))\eta s\leq\left(1-\eta\left(\lambda-2\frac{19}{28}\lambda\right)\right)\frac{\eta\lambda}{3L}=\frac{\eta\lambda(5\eta\lambda+14)}{42L}.$
By taking (4.16) and (4.17) into (4.14), one can get
$\mathbb{E}_{\xi}\left[{}\left\|G(x+\sigma(x)\xi)-x^{*}\right\|^{2}\right]\leq\left(1-\frac{\eta\lambda}{100}\right)\left\|x-x^{*}\right\|^{2}+\frac{\eta\lambda(5\eta\lambda+14)}{42L}(f^{*}-f_{\textsc{lb}}),$
which is (4.6), and this completes the proof. ∎
We are ready to give the proof of Theorem 3.1.
Proof of Theorem 3.1.
In accordance with Lemma 4.4, $G(x)=x-\eta\nabla f(x)$ is the gradient descent map. Denote $\omega_{t}:={\rm SG}(x_{t})-\nabla f(x_{t})$ as the random noise so that by 3.1 we know that
(4.18) $\mathbb{E}\left[{}\omega_{t}\right]=0\quad\mbox{and}\quad\mathbb{E}\left[{}\left\|\omega_{t}\right\|^{2}\right]\leq r^{2}.$
By letting $y_{t}:=x_{t}-\eta\nabla f(x_{t})$ and $\sigma(x):=\sqrt{\eta s\left(f(x)-f_{\textsc{lb}}\right)}$, one can rewrite the iterations in Algorithm 1 as follows
$\left\{\begin{array}[]{l}y_{t}=x_{t}-\eta\nabla f(x_{t})=G(x_{t}),\\ x_{t+\frac{1}{2}}=x_{t}-\eta{\rm SG}(x_{t})=y_{t}-\eta\omega_{t},\\ x_{t+1}=x_{t+\frac{1}{2}}-\sigma_{t}\xi_{t}=x_{t+\frac{1}{2}}-\sigma(x_{t+\frac{1}{2}})\xi_{t},\end{array}\right.\quad 0\leq t\leq T,$
which can be reorganized as
(4.19) $\left\{\begin{array}[]{l}x_{t+\frac{1}{2}}=y_{t}-\eta\omega_{t},\\ x_{t+1}=x_{t+\frac{1}{2}}-\sigma(x_{t+\frac{1}{2}})\xi_{t},\\ y_{t+1}=G(x_{t+1}),\end{array}\right.\quad\quad 0\leq t\leq T.$
Note that the $\sigma$-fields $\mathcal{F}_{t}=\sigma\{\xi_{0},\cdots,\xi_{t-1},\omega_{0},\cdots,\omega_{t-1}\}$ and $\mathcal{F}_{t+\frac{1}{2}}=\sigma\{\xi_{0},\cdots,\xi_{t-1},\omega_{0},\cdots,\omega_{t-1},\omega_{t}\}$ constitute a generated filtration. We know from Lemma 2.1 that $x^{*}$ is the unique global minima of $f$, so that by (4.19) we have
[…]
Then by using Lemma 4.4, (3.2) and (4.18) we can get
[…]
$\leq\left(1-\frac{\eta\lambda}{100}\right)\left\|y_{t}-x^{*}\right\|^{2}+\eta^{2}r^{2}+\frac{\eta\lambda(5\eta\lambda+14)}{42L}(f^{*}-f_{\textsc{lb}})$
$=\left(1-\frac{\eta\lambda}{100}\right)\left\|y_{t}-x^{*}\right\|^{2}+\eta\lambda b.$
Consequently,
[…]
By taking expectation on both sides of (4.20) we can get
(4.21) $\mathbb{E}\left[{}\left\|y_{t+1}-x^{*}\right\|^{2}-100b\right]\leq\left(1-\frac{\eta\lambda}{100}\right)\mathbb{E}\left[{}\left\|y_{t}-x^{*}\right\|^{2}-100b\right].$
In case $b=0$, the inequality (4.21) turns to $\mathbb{E}\left[{}\left\|y_{T}-x^{*}\right\|^{2}\right]\leq\left(1-\frac{\eta\lambda}{100}\right)^{T}\left\|y_{0}-x^{*}\right\|^{2}$. Thus, for any $\varepsilon>0$, Markov’s inequality implies
$\mathbb{P}\left[{}\left\|y_{T}-x^{*}\right\|^{2}>\varepsilon\right]\leq\frac{1}{\varepsilon}\left(1-\frac{\eta\lambda}{100}\right)^{T}\left\|y_{0}-x^{*}\right\|^{2}.$
Consequently, for any $\zeta\in(0,1)$, one has $\left\|y_{T}-x^{*}\right\|^{2}\leq\varepsilon$ with the probability at least being $1-\zeta$, as long as
$T\geq\frac{100}{\eta\lambda}\ln\frac{\left\|y_{0}-x^{*}\right\|^{2}}{\zeta\varepsilon}.$
On the other hand, if $b>0$, by $\eta\lambda\leq\frac{16}{25}$, we can use (4.20) by considering the stochastic process $\{\left\|y_{t}-x^{*}\right\|^{2}-100b\}_{t\in\mathds{N}}$. More specifically, by applying Lemma 4.3 with $\theta=1-\frac{\eta\lambda}{100}$ and $\ell=100b$ to this stochastic process one can get that for any given nonnegative integer $T$,
$\mathbb{P}\left[{}\exists t\leq T,\left\|y_{t}-x^{*}\right\|^{2}<200b\right]\geq 1-\left(\frac{200-{\eta\lambda}}{200}\right)^{T}\frac{B}{100b}$
with
$B=\left(\left\|y_{0}-x^{*}\right\|^{2}-100b\right)\mathbb{P}\left[{}\left\|y_{0}-x^{*}\right\|^{2}\geq 200b\right]\leq\left\|y_{0}-x^{*}\right\|^{2}.$
By taking $T\geq\frac{200}{\eta\lambda}\ln\frac{\left\|y_{0}-x^{*}\right\|^{2}}{100b\zeta}$ one obtains that
$1-\left(1-\frac{\eta\lambda}{200}\right)^{T}\frac{B}{100b}\geq 1-\left(1-\frac{\eta\lambda}{200}\right)^{T}\frac{\left\|y_{0}-x^{*}\right\|^{2}}{100b}\geq 1-\zeta,$
Furthermore, note that $y_{t}=x_{t}-\eta\nabla f(x_{t})$ and $\eta L\leq 2/5$. One has from (2.1) that
(4.22) $\left\|y_{t}-x^{*}\right\|\geq\left\|x_{t}-x^{*}\right\|-\eta\left\|\nabla f(x_{t})\right\|\geq\left\|x_{t}-x^{*}\right\|-\eta L\left\|x_{t}-x^{*}\right\|\geq(1-\eta L)\left\|x_{t}-x^{*}\right\|.$
Then, from (2.5) of Lemma 2.1 we have
(4.23) $f(x_{t})-f^{*}\leq\frac{L}{2}\left\|x_{t}-x^{*}\right\|^{2}\leq\frac{L}{2(1-\eta L)^{2}}\left\|y_{t}-x^{*}\right\|^{2}.$
According to Theorem 3.1, with probability at least $1-\zeta$, there exists $t\leq T$ such that $\left\|y_{t}-x^{*}\right\|^{2}\leq 200b$. Thus, by (4.23) we have (3.3) holds, and it, together with Lemma 2.1, implies (3.4). Similar to (4.22), we can get
(4.24) $\left\|y_{t}-x^{*}\right\|\leq\left\|x_{t}-x^{*}\right\|+\eta\left\|\nabla f(x_{t})\right\|\leq\left\|x_{t}-x^{*}\right\|+\eta L\left\|x_{t}-x^{*}\right\|\leq(1+\eta L)\left\|x_{t}-x^{*}\right\|$
. Then by (3.4) we know that
$\left\|y_{t_{*}}-x^{*}\right\|^{2}\leq(1+\eta L)^{2}\frac{200Lb}{(1-\eta L)^{2}(\alpha-\beta)},$
which implies (3.5), and this completes the proof. ∎
4.3. Proof of Theorem 3.2
Proof of Theorem 3.2.
Define
$b^{\nu}=\frac{\eta r^{2}}{\lambda}+\frac{5\eta\lambda+14}{42L}(f^{*}-f_{\textsc{lb}}^{\nu}),\quad\forall\nu>0.$
Meanwhile, define the events
$\mathfrak{E}_{\nu}=\{f^{*}-f_{\textsc{lb}}^{i}>\varepsilon,\forall i<\nu\}\cap\{\forall i<\nu,\exists t_{i}\leq T_{2},\left\|y_{t_{i}}^{i}-x^{*}\right\|^{2}\leq 200b^{i}\},\quad\forall\nu\geq 0,$
so that $\mathfrak{E}_{0}$ is defined as the total probability space. Obviously, $\mathfrak{E}_{\nu}\supset\mathfrak{E}_{\nu+1}$. Moreover, if $\mathfrak{E}_{\nu}$ holds for a certain $\nu>0$, by (4.23) we know that
(4.25) $f^{*}-f_{\textsc{lb}}^{\nu-1}>\varepsilon\quad\mbox{and}\quad f(x_{\min}^{\nu-1})-f^{*}\leq\frac{L}{2(1-\eta L)^{2}}\min_{t}\left\|y^{\nu-1}_{t}-x^{*}\right\|^{2}\leq\frac{100Lb^{\nu-1}}{(1-\eta L)^{2}}.$
Consequently, if $\mathfrak{E}_{\nu}$ holds, by using the facts that $b^{\varepsilon}=\frac{\eta r^{2}}{\lambda}+\frac{5\eta\lambda+14}{42L}\varepsilon$ and $\gamma=\frac{(1-\eta L)^{2}\varepsilon}{(1-\eta L)^{2}\varepsilon+100Lb^{\varepsilon}}$, one has
(4.26) $f^{*}-f_{\textsc{lb}}^{\nu}$ $=(1-\gamma)\left(f^{*}-f_{\textsc{lb}}^{\nu-1}\right)-\gamma\left(f(x_{\min}^{\nu-1})-f^{*}\right)$
$\geq(1-\gamma)\left(f^{*}-f_{\textsc{lb}}^{\nu-1}\right)-\gamma\frac{100L}{(1-\eta L)^{2}}b^{\nu-1}$
$=\frac{100L}{(1-\eta L)^{2}\varepsilon+100L}\left(\left(f^{*}-f_{\textsc{lb}}^{\nu-1}\right)b^{\varepsilon}-\varepsilon b^{\nu-1}\right)$
$=\frac{100L}{(1-\eta L)^{2}\varepsilon+100L}\frac{\eta r^{2}}{\lambda}\left({f^{*}-f_{\textsc{lb}}^{\nu-1}}-\varepsilon\right)>0.$
Based on (4.26), we define the events
$\mathfrak{A}_{\nu}=\left\{0\leq f^{*}-f_{\textsc{lb}}^{\nu}\leq\varepsilon\right\}\quad\mbox{and}\quad\mathfrak{B}_{\nu}=\left\{\left\|y_{t}^{\nu}-x^{*}\right\|^{2}>200b^{\nu},\forall t\leq T_{2}\right\},\quad\forall\nu\geq 0.$
According to the above discussions we know that $\mathfrak{E}_{\nu}\backslash\mathfrak{E}_{\nu+1}=\mathfrak{E}_{\nu}\cap(\mathfrak{A}_{\nu}\cup\mathfrak{B}_{\nu})$. Note that $b^{\nu}>b^{\varepsilon}$ is equivalent to $f^{*}-f_{\textsc{lb}}^{\nu}>\varepsilon$. Consequently, by defining
$\mathfrak{B}_{\nu}^{\prime}=\begin{cases}\mathfrak{B}_{\nu},&\nu=0,\\ \left\{\forall t\leq T_{2},\left\|y_{t}^{\nu}-x^{*}\right\|^{2}>200\cdot\max\{b^{\nu},b^{\varepsilon}\}\right\},&\forall\nu>0,\end{cases}$
one has $\mathfrak{A}_{\nu}\cup\mathfrak{B}_{\nu}=\mathfrak{A}_{\nu}\cup\mathfrak{B}_{\nu}^{\prime},\forall\nu>0$. Therefore,
(4.27) $\mathbb{P}\left[{}\mathfrak{E}_{\nu}\right]$ $=\mathbb{P}\left[{}\mathfrak{E}_{\nu+1}\right]+\mathbb{P}\left[{}\mathfrak{E}_{\nu}\cap(\mathfrak{A}_{\nu}\cup\mathfrak{B}_{\nu}^{\prime})\right]$
$=\mathbb{P}\left[{}\mathfrak{E}_{\nu+1}\right]+\mathbb{P}\left[{}\mathfrak{E}_{\nu}\cap\mathfrak{A}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime c}\right]+\mathbb{P}\left[{}\mathfrak{E}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime}\right]\quad\forall\nu\geq 0.$
In the following, we prove
(4.28) $\mathbb{P}\left[{}\mathfrak{B}_{\nu}^{\prime}\mid\mathfrak{E}_{\nu}\right]\leq\zeta^{\prime},\quad\forall\nu\geq 0.$
According to Theorem 3.1 and the bound of $T_{1}$ given in (3.8), we know that (4.28) holds for $\nu=0$. For the case that $\nu>0$, we know that $\mathfrak{E}_{\nu}$ happens implies that (4.26) holds. Moreover, it holds that
$\mathbb{P}\left[{}\mathfrak{B}_{\nu}^{\prime}\mid\mathfrak{E}_{\nu}\right]\leq\mathbb{P}\left[{}\left\{\forall t\leq T_{2},\left\|y_{t}^{\nu}-x^{*}\right\|^{2}>200\cdot b^{\nu}\right\}\mid\mathfrak{E}_{\nu}\right].$
Thus, if $b^{\nu}\geq b^{\varepsilon}$, by Theorem 3.1 we know that (4.28) holds if
(4.29) $T_{2}\geq\frac{200}{\eta\lambda}\ln\frac{\left\|y_{0}^{\nu}-x^{*}\right\|^{2}}{100b^{\varepsilon}\zeta^{\prime}}\geq\frac{200}{\eta\lambda}\ln\frac{\left\|y_{0}^{\nu}-x^{*}\right\|^{2}}{100b^{\nu}\zeta^{\prime}}.$
On the other hand, suppose that $\nu>0$ and $b^{\nu}<b^{\varepsilon}$. Similar to the proof of Theorem 3.1, we define
(4.30) $G^{\nu}_{t}=\left\|y_{t}^{\nu}-x^{*}\right\|^{2}-100b^{\nu}\geq-100b^{\nu}.$
Therefore, with $\mathfrak{E}_{\nu}$ holds, the generated filtration is given by
$\mathcal{F}^{\nu}_{t}:=\sigma\{\xi^{\nu}_{0},\cdots,\xi^{\nu}_{t-1},\omega^{\nu}_{0},\cdots,\omega^{\nu}_{t-1}\},\quad\forall\nu>0,0\leq t\leq T_{2};$
We know that (cf. the derivation of eq. 4.20) that,
$\mathbb{E}\left[{}G^{\nu}_{t+1}\mid\mathcal{F}_{t}\right]\leq\left(1-\frac{\eta\lambda}{100}\right)G^{\nu}_{t},\quad\forall 0\leq t\leq T_{2}-1,\ \forall\nu>0\mbox{ with }\mathfrak{E}_{\nu}\mbox{ holds}.$
Taking $\theta=1-\frac{\eta\lambda}{100}$, $b=100b^{\nu}$ and $\ell=100b^{\epsilon}$ in Lemma 4.3, one can get
(4.31) $\mathbb{P}\left[{}\exists t\leq T_{2},G^{\nu}_{t}<100b^{\varepsilon}\right]\geq 1-\left(1-\frac{\eta\lambda}{100}\frac{b^{\varepsilon}}{b^{\nu}+b^{\varepsilon}}\right)^{T}\frac{B}{100b^{\varepsilon}}\quad\mbox{with}\quad B:=\mathbb{E}\left[{}G^{\nu}_{0}\mathds{1}\{G^{\nu}_{0}\geq 100b^{\epsilon}\}\mid\mathfrak{E}_{\nu}\right]\geq 0.$
From the (4.30) we know that $B\leq\left\|y_{0}^{\nu}-x^{*}\right\|^{2}$. Meanwhile, $b^{\nu}<b^{\varepsilon}$ implies $\frac{b^{\varepsilon}}{b^{\nu}+b^{\varepsilon}}\geq\frac{1}{2}$. Hence, (4.31) implies
$\mathbb{P}\left[{}\exists t\leq T_{2},G^{\nu}_{t}<100b^{\varepsilon}\right]\geq 1-\left(1-\frac{\eta\lambda}{200}\right)^{T}\frac{\left\|y_{0}^{\nu}-x^{*}\right\|^{2}}{100b^{\varepsilon}}.$
Therefore, (4.28) holds if one has
(4.32) $T_{2}\geq\frac{200}{\eta\lambda}\ln\frac{\left\|y_{0}^{\nu}-x^{*}\right\|^{2}}{100b^{\varepsilon}\zeta^{\prime}},$
which, together with (4.29), implies that (4.32) is sufficient for (4.28).
Note that whenever $\mathfrak{E}_{\nu}$ holds, $x_{0}^{\nu}=x_{\min}^{\nu-1}$. Then, by using (4.24) and Lemma 2.1 one can obtain
$\left\|y_{0}^{\nu}-x^{*}\right\|^{2}\leq(1+\eta L)^{2}\|x_{0}^{\nu}-x^{*}\|^{2}\leq\frac{2(1+\eta L)^{2}}{\alpha-\beta}\left(f(x_{\min}^{\nu-1})-f^{*}\right)$
Taking (4.25) into account, one gets
$\left\|y_{0}^{\nu}-x^{*}\right\|^{2}\leq\left(\frac{1+\eta L}{1-\eta L}\right)^{2}\frac{200Lb^{\nu-1}}{\alpha-\beta}.$
Moreover, since $f(x_{\min}^{\nu-1})\geq f^{*}$ and $f_{\textsc{lb}}^{\nu}=(1-\gamma)f_{\textsc{lb}}^{\nu-1}+\gamma f\left(x_{\min}^{\nu-1}\right)$, one has
(4.33) $f^{*}-f_{\textsc{lb}}^{\nu}=(1-\gamma)\left(f^{*}-f_{\textsc{lb}}^{\nu-1}\right)-\gamma\left(f(x_{\min}^{\nu-1})-f^{*}\right)\leq f^{*}-f_{\textsc{lb}}^{\nu-1}.$
Thus, by $\mathfrak{E}_{0}\supset\cdots\supset\mathfrak{E}_{\nu-1}\supset\mathfrak{E}_{\nu}$ one has $b^{\nu-1}\leq b^{0}$, so that
$\frac{200}{\eta\lambda}\ln\frac{\left\|y_{0}^{\nu}-x^{*}\right\|^{2}}{100b^{\varepsilon}\zeta^{\prime}}\leq\frac{200}{\eta\lambda}\ln\left(\left(\frac{1+\eta L}{1-\eta L}\right)^{2}\frac{2Lb^{0}}{(\alpha-\beta)b^{\varepsilon}\zeta^{\prime}}\right)\leq T_{2}.$
Therefore, (4.28) always holds.
Based on (4.28), by using (4.27) one can get
(4.34) $\mathbb{P}\left[{}\mathfrak{E}_{\nu}\right]-\mathbb{P}\left[{}\mathfrak{E}_{\nu+1}\right]=\mathbb{P}\left[{}\mathfrak{E}_{\nu}\cap\mathfrak{A}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime c}\right]+\mathbb{P}\left[{}\mathfrak{E}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime}\right]\leq\mathbb{P}\left[{}\mathfrak{E}_{\nu}\cap\mathfrak{A}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime c}\right]+\zeta^{\prime}.$
Summing (4.34) up with $\nu$ from 00 to $N$ and using the fact that $\mathbb{P}\left[{}\mathfrak{E}_{0}\right]=1$, one has
$\mathbb{P}\left[{}\mathfrak{E}_{N+1}\right]+\sum_{\nu=0}^{N}\mathbb{P}\left[{}\mathfrak{E}_{\nu}\cap\mathfrak{A}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime c}\right]\geq 1-(N+1)\zeta^{\prime}=1-\zeta$
Note that $\mathfrak{E}_{\nu}\cap\mathfrak{A}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime c}\subseteq\mathfrak{E}_{\nu}\backslash\mathfrak{E}_{\nu+1}$ for all $\nu\geq 0$, so that the sets $\mathfrak{E}_{\nu}\cap\mathfrak{A}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime c}$ are separated for all $\nu\geq 0$. Furthermore, each event $\mathfrak{E}_{\nu}\cap\mathfrak{A}_{\nu}\cap\mathfrak{B}_{\nu}^{\prime c}\subseteq\mathfrak{E}_{\nu}\backslash\mathfrak{E}_{\nu+1}$ means that there exists $t\leq T_{2}$ (or $t\leq T_{1}$ when $\nu=0$) such that $\left\|y_{t}^{\nu}-x^{*}\right\|^{2}\leq 200b^{\varepsilon}$, which yields from assentation (b) of Theorem 3.1 that $f\left(x_{\min}^{\nu}\right)-f^{*}\leq\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}$. Due to the monotonicity fact that $f(x_{\min}^{N})\leq f(x_{\min}^{\nu})$, we obtain $f\left(x_{\min}^{N}\right)-f^{*}\leq\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}$. Therefore, the last statement that should be proved is
(4.35) $\mathfrak{E}_{N+1}\Rightarrow f\left(x_{\min}^{N}\right)-f^{*}\leq\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}},$
so that by Lemma 2.1 one has $\left\|x_{\min}^{N}-x^{*}\right\|^{2}\leq\frac{200Lb^{\varepsilon}}{(1-\eta L)^{2}(\alpha-\beta)}$.
Suppose on the contrary that (4.35) is not true. Since $f(x_{\min}^{N})\leq f(x_{\min}^{\nu})$ for all $\nu\geq 0$, we know that $f\left(x_{\min}^{\nu}\right)-f^{*}\geq\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}$ holds for all $\nu=1,\ldots,N$. Therefore, from (4.33) we know that
$f^{*}-f_{\textsc{lb}}^{\nu}=(1-\gamma)\left(f^{*}-f_{\textsc{lb}}^{\nu-1}\right)-\gamma\left(f(x_{\min}^{\nu-1})-f^{*}\right)\leq(1-\gamma)\left(f^{*}-f_{\textsc{lb}}^{\nu-1}\right)-\gamma\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}},$
which is equivalent to
$\left(f^{*}-f_{\textsc{lb}}^{\nu}+\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}\right)\leq(1-\gamma)\left(f^{*}-f_{\textsc{lb}}^{\nu-1}+\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}\right).$
Therefore, as $\mathfrak{E}_{N+1}$ implies $f^{*}-f_{\textsc{lb}}^{N}>\varepsilon$ via (4.25),
$\varepsilon+\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}\leq(1-\gamma)^{N}\left(f^{*}-f_{\textsc{lb}}^{0}+\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}\right).$
Then by using $\gamma=\frac{(1-\eta L)^{2}\varepsilon}{(1-\eta L)^{2}\varepsilon+100Lb^{\varepsilon}}=\frac{\varepsilon}{\varepsilon+\frac{100Lb^{\varepsilon}}{(1-\eta L)^{2}}}$ given in (3.7), we have
$\frac{\varepsilon}{\gamma}\leq(1-\gamma)^{N}\left(f^{*}-f_{\textsc{lb}}^{0}-\varepsilon+\frac{\varepsilon}{\gamma}\right)\Leftrightarrow(1-\gamma)^{-N}\leq(1-\gamma)+\gamma\cdot\frac{f^{*}-f_{\textsc{lb}}^{0}}{\varepsilon},$
which contradicts to $N\geq\frac{1}{\gamma}\ln\left((1-\gamma)+\gamma\cdot\frac{f^{*}-f_{\textsc{lb}}^{0}}{\varepsilon}\right)$. Therefore, (4.35) holds.
Overall, we have proved that, with probability at least $1-\zeta$, one has $\left\|x_{\min}^{N}-x^{*}\right\|^{2}\leq\frac{200Lb^{\varepsilon}}{(1-\eta L)^{2}(\alpha-\beta)}$, and this completes the proof of the theorem. ∎
5. Numerical Experiments
In this section, we conduct numerical experiments to test the performance of the proposed GND (Algorithm 1) and DL-GND (Algorithm 2) algorithms. By sampling 10,000 initial points, we run the algorithms and evaluate their performance using the mean squared error (MSE) and the non-convergence probability (N-CP). These two measures are defined by
(5.1) $\text{MSE}(t)=\mathbb{E}(\|x_{t}-x^{*}\|^{2}),\quad\text{N-CP}(t)=\mathbb{P}(\|x_{t}-x^{*}\|>10^{-3}),$
where $x_{t}$ is the random vector of $t$-th iteration and $x^{*}$ is the global minima. The MSE quantifies convergence in the state space, while the N-CP quantifies convergence in the probability space.
5.1. Minimizing a uni-variate nearly convex function
Recall that the objective function defined in Example 2.1 is given by
(5.2) $f_{n,k}(x):=\frac{1}{2}x^{2}-\left(1+\frac{1}{k}\right)\int_{0}^{x}t(\sin t)^{2n}\mathrm{d}t,\quad x\in,$
where $n,k\in\mathbb{N}^{+}$. Here, we consider two cases: $n=7,k=1$ and $n=112,k=2$. Notably, we can prove that $f_{112,2}$ is $\left(\frac{21}{25},1\right)$-nearly convex. However, it remains undetermined whether $f_{7,1}$ is nearly convex as the theoretical bound in (2.2) may not be satisfied.
(a) $n=112,k=2$
(b) $n=7,k=1$
Figure 1. The MSE and N-CP versus iterations of the GD, GND, and DL-GND algorithms are evaluated by minimizing $f_{7,1}(x)$ and $f_{112,2}(x)$ given by Equation 5.2.
We run the GND and DL-GND algorithms choosing 10,000 random initial points from the interval $[-10,10]$. In this experiment, we test the full gradient case, that is, ${\rm SG}(x)=\nabla f(x)$ and $r=0$. We set $\eta=0.4,s=0.5$ for the case $n=7,k=1$ and $\eta=0.1,s=0.2$ for the case $n=112,k=2$. The difference in parameter settings arises as the oscillation of $f_{7,1}(x)$ is more severe, requiring larger values for $\eta$ and $s$. In GND algorithm, we set $f_{\textsc{lb}}=0$. For the DL-GND case, we set $f_{\textsc{lb}}=-1$ and take $\gamma=0.5,T_{1}=40,T_{2}=10,N=30$.
The numerical behaviors of GND and DL-GND are given in Figure 1 in terms of the MSE and the N-CP. Both GND and DL-GND (Algorithms 1 and 2) algorithms are observed to converge linearly to the global optimal point in both state space and probability space, while most trajectories of GD fail to converge and get trapped in local minima. Due to the additional estimation of the lower bound of the objective function value, DL-GND requires more iterations than GND, but achieves the similar accuracy.
These results demonstrate our theoretical findings (Theorems 3.1 and 3.2). In addition, although we can not prove the nearly convexity of $f_{7,1}(x)$, the numerical results show the effectiveness of GND and DL-GND algorithms, suggesting potential for future research on identifying weaker conditions for global convergence.
5.2. Minimizing Rastrigin functions
In this subsection, we test the performance of GND and DL-GND algorithms for minimizing the Rastrigin function [16]:
(5.3) $f(\bm{x}):=a\left(d-\sum_{i=1}^{d}\cos\left(bx^{i}\right)\right)+c\sum_{i=1}^{d}(x^{i})^{2},\quad\bm{x}=(x^{1},\ldots,x^{d})^{T}\in^{d}.$
Here, we set $a=b=1$ and vary the dimension $d$ and the coefficient $c$.
5.2.1. Two dimensional case
We perform numerical experiments to minimize the Rastrigin function $f(x)$ defined in (5.3) with $d=2$, focusing on the cases $c=0.01$ and $c=0.05$. The proposed GND and DL-GND algorithms are applied to $f(x)$, and their performance is evaluated using MSE and N-CP as defined in (5.1). We adopt the following common setup for all experiments: Each consists of 10,000 independent and identically distributed (i.i.d.) trajectories to estimate the MSE and the N-CP. Each trajectory originates from a uniform distribution over the region $[-20,20]^{2}$. Recall that $x_{t}$ is the random vector of $t$-th iteration. The experiments are divided into the following two parts.
Part I: Default performance evaluation.
In this part, we test the GND and DL-GND algorithms with their default optimal parameters, given in Table 1. It is noted that in GND algorithm, $f_{\textsc{lb}}=f^{*}=0$.
Table 1. Default parameter configurations for minimizing rastrigin functions.
Dimension $c$ Algorithm $\eta$ $s$ $f_{\textsc{lb}}$ $f_{\textsc{lb}}^{0}$ $\gamma$ $T_{1}$ $T_{2}$
$d=2$ $c=0.05$ GND 1.5 2.0 0 - - - -
DL-GND 1.5 1.5 - -20 0.3 100 10
$c=0.01$ GND 1.5 4.0 0 - - - -
DL-GND 1.5 3.0 - -20 0.03 100 10
$d=10$ $c=0.05$ GND 1.5 1.5 0 - - - -
DL-GND 1.5 1.4 - -20 0.025 100 10
$c=0.03$ GND 1.5 2.5 0 - - - -
DL-GND 1.5 1.5 - -20 0.0035 100 10
(a) $c=0.05,d=2$
(b) $c=0.01,d=2$
Figure 2. The convergence behavior for minimizing the two dimensional Rastrigin function (5.3) with $c=0.05$ and $0.01$. The first and the second rows are the MSE and N-CP versus iterations, respectively.
Figure 3. One trajectory for minimizing the two dimensional Rastrigin function (5.3) with $c=0.01$ and $0.05$ using the DL-GND algorithms. The convergence behavior of function value (Left) and the estimated lower bound of $f_{\textsc{lb}}^{\nu}$ (Right).
The numerical results of GND and DL-GND, shown in Figure 2, are evaluated based on the metrics MSE and N-CP. The results show that both GND and DL-GND achieve linear convergence to the global minimum for $c=0.05$ and $c=0.01$, in both the state space and the probability space. However, the DL-GND algorithm converges 4–8 times more slowly than the GND algorithm for both objectives. This slower convergence is considered acceptable, given that the initial gap $f^{*}-f_{\textsc{lb}}^{0}=20$ in DL-GND is significantly larger compared to the zero gap in GND.
Figure 3 illustrates a single trajectory of DL-GND, showing both the function values ($f(x_{t})$) and the adaptive lower bounds ($f_{\textsc{lb}}^{\nu}$). The result reveal that the lower bound estimate $f_{\textsc{lb}}^{\nu}$ converges linearly to $f^{*}$, while the function value $f(x_{t})$ also progressively converges to $f^{*}$, accompanied by noise-induced oscillations.
The experiments in the first part demonstrate that both GND and DL-GND efficiently achieve global convergence. Notably, for DL-GND, even in the absence of prior knowledge of the exact value of $f^{*}$, the adaptive strategy guarantees the linear convergence of $f_{\textsc{lb}}^{\nu}$ to $f^{*}$.
Part II: Parameter sensitivity analysis.
In this part, we study the effect of different parameter choices on the convergence of the algorithms. The GND and DL-GND algorithms are applied to optimize $f(x)$ in (5.3) with $d=2$ and $c=0.01$. The GND algorithm involves three hyperparameters: the learning rate or step size $\eta$, the noise coefficient $s$, and the lower bound of the function value $f_{\textsc{lb}}$. We compare the results obtained by varying the hyperparameter $\eta,s$ and $f_{\textsc{lb}}$, as shown in Figures 4 and 5. For the DL-GND algorithm, the key parameter influencing the convergence is $\gamma$, which controls the lower bound estimation of the function value. The corresponding results are presented in Figure 6.
Figure 4. Minimizing the two dimensional Rastrigin function (5.3) with $c=0.01$, using the GND algorithms. We fix $f_{\textsc{lb}}=f^{*}=0$ and vary the two parameters $\eta$ and $s$.
Figure 5. Minimizing the two-dimensional Rastrigin function (5.3) with $c=0.01$, using the GND algorithms. We fix $\eta=1.5,s=3.0$ and vary $f_{\textsc{lb}}$ from $-1$ to $-10^{-10}$.
Figure 6. Minimizing the two-dimensional Rastrigin function (5.3) with $c=0.01$, using the DL-GND algorithms. We fix $\eta=1.5,s=3.0,f_{\textsc{lb}}^{0}=-20,T_{1}=100,T_{2}=10$ and vary $\gamma$ from $0.02$ to $0.04$.
In Figure 4, the GND algorithm is applied with the lower bound fixed at $f_{\textsc{lb}}=f^{*}=0$, while varying the values of $\eta$ and $s$. The hyperparameters are selected from the set
$(\eta,s)\in\{0.5,1.0,1.5\}\times\{1.0,2.5,5.0\}.$
The estimated MSE and N-CP for each combination of $(\eta,s)$ are summarized in Figure 4. The results indicate that GND converges linearly to the global minimum for all choices of $\eta$ when $s\geq 2.5$. However, when the noise coefficient $s$ is too small, the algorithm fails to converge to the global minimum, suggesting that insufficient noise makes it harder for the trajectory to escape local minima. When $s$ is chosen suitably large, the choice of $\eta$ and $s$ affects the convergence speed. Larger $\eta$ and $s$ appear to converge faster, but the converging trajectories become noisier. To achieve satisfactory performance, $\eta$ and $s$ should not be too large.
In Figure 5, the GND algorithm is applied with fixed $\eta=1.5,s=3.0$, while varying $f_{\textsc{lb}}$ from $-1$ to $-10^{-10}$. The results indicate that $f_{\textsc{lb}}$ plays a critical role in the global convergence of GND. Specifically, convergence to within 0.001 of $x^{*}$ is not achieved until $f^{*}-f_{\textsc{lb}}\leq 10^{-8}$.
In Figure 6, we apply the DL-GND algorithm with fixed parameters $\eta=1.5,s=3.0,f_{\textsc{lb}}^{0}=-20$ and varying $\gamma$ from $0.02$ to $0.04$. The number of iterations is set to $T_{1}=100$ for the initial stage and $T_{2}=10$ for the inner loop. The results indicate that $\gamma$ significantly affect the convergence of the DL-GND algorithm. When $\gamma$ is too small, the convergence speed is excessively slow. Conversely, when $\gamma$ is too large, some trajectories fail to converge as the lower bound estimations $f_{\textsc{lb}}^{\nu}$ grow too quickly, occasionally surpassing $f^{*}$ in some outer loops. Therefore, selecting an appropriate $\gamma$ is essential to achieve efficient convergence in DL-GND.
The experiments in the second part demonstrate that, with appropriate choices of hyperparameters $\eta$ and $s$, and an accurate lower bound estimation of $f$, the GND algorithm can efficiently converge to a neighborhood of the global minimum for certain nonconvex functions. While the DL-GND algorithm eliminates the need for an accurate lower bound estimation, it requires a properly chosen $\gamma$ to efficiently update the lower bound estimation while ensuring the constraint $f_{\textsc{lb}}^{\nu}\leq f^{*}$ is satisfied. In practical implementations, these hyperparameters can be tuned to achieve optimal performance.
5.2.2. Ten dimensional case
We perform experiments to minimize the 10-dimensional Rastrigin function $f(x)$ defined in eq. 5.3 with $c=0.05$ and $c=0.03$. The proposed GND and DL-GND algorithms are applied to $f(x)$, and their performance is evaluated using mean squared error (MSE) and non-convergence probability (N-CP) as defined in (5.1).
For the GND algorithm, we set $f_{\textsc{lb}}=f^{*}=0$, while for the DL-GND algorithm, we set $f_{\textsc{lb}}^{0}=-20$. All other hyperparameters, including $\eta,s,\gamma$, are treated as tunable parameters, with the detailed parameter configuration provided in Table 1. The remaining experimental setup is analogous to the 2-dimensional case. The results, evaluated in terms of MSE and N-CP, are displayed in Figure 7.
From Figure 7, we observe that both the GND and DL-GND algorithm successfully converge to the global minimum in both settings. In the 10-dimensional space, the convergence speed is nearly 10 times slower compared to the 2-dimensional case. This slowdown is expected, as the number of local minima and saddle points grows exponentially with the dimension. These results highlight the robustness and effectiveness of the proposed GND and DL-GND algorithms in handling high-dimensional nonconvex optimization problems.
(a) $c=0.05,d=10$
(b) $c=0.03,d=10$
Figure 7. The convergence behavior for minimizing the ten dimensional Rastrigin function (5.3) with $c=0.05$ and $0.03$. The first and the second rows are the MSE and N-CP versus iterations, respectively.
6. Conclusions
This study introduces a SGD method with an adaptive Gaussian noise mechanism and proves its global convergence for minimizing nearly convex functions. The Gaussian noise added at each step is determined by the difference between the current and optimal values of the objective function. Additionally, we propose a double-loop iterative scheme that estimates the lower bound of the objective function without requiring its exact optimal value. Empirical results validate the effectiveness of the proposed algorithms. Future work could explore extensions to other stochastic methods and broader classes of objective functions to further generalize these findings.
Appendix A Appendix
A.1. Proof of Lemma 4.2
Proof.
Define the function $\vartheta(x):=\varepsilon(x+x^{*})$, so that $\vartheta(0)=0$ and $|\vartheta(x)|\leq\frac{l}{2}\left\|x\right\|^{2}$. Moreover, it holds that
$\nabla\varepsilon(x-\delta\xi)=\nabla\varepsilon(x-\delta\xi-x^{*}+x^{*})=\nabla\vartheta(x-x^{*}-\delta\xi)=\nabla\vartheta(v-\delta\xi)\quad\mbox{with}\quad v:=x-x^{*}.$
Note that the right-hand side of (4.2) is only related to the length of $v=x-x^{*}$. Therefore, without loss of generality, we can assume $v=\left\|v\right\|e_{1}$, where $e_{1}$ is the unit vector: $e_{1}=(1,0,\cdots,0)^{\top}$. Consequently,
$\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\varepsilon(x-\delta\xi),x-x^{*}\rangle\right]\right\rvert=\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\vartheta(v-\delta\xi),v\rangle\right]\right\rvert=\left\lvert\mathbb{E}_{\xi}\left[{}\partial_{v_{1}}\vartheta(v-\delta\xi)\right]\right\rvert\cdot\left\|v\right\|.$
By denoting $\tilde{\vartheta}(\xi)=\vartheta(v-\delta\xi)$ one has $\partial_{\xi_{1}}\tilde{\vartheta}(\xi)=-\delta\partial_{v_{1}}\vartheta(v-\delta\xi)$. Note that $\rho(\xi)=\frac{d^{\frac{d}{2}}}{(2\pi)^{\frac{d}{2}}}\exp\left(-\frac{d\left\|\xi\right\|^{2}}{2}\right)$ so that $\nabla\rho(\xi)=-d\rho(\xi)\xi$. Therefore, due to partial integration,
$\mathbb{E}_{\xi}\left[{}\partial_{\xi_{1}}\tilde{\vartheta}(\xi)\right]=\int_{{}^{d}}\partial_{\xi_{1}}\tilde{\vartheta}(\xi)\rho(\xi)\mathrm{d}\xi=-\int_{{}^{d}}\tilde{\vartheta}(\xi)\partial_{\xi_{1}}\rho(\xi)\mathrm{d}\xi=d\int_{{}^{d}}\tilde{\vartheta}(\xi)\rho(\xi)\xi_{1}\mathrm{d}\xi.$
Consequently,
$\left\lvert\mathbb{E}_{\xi}\left[{}\partial_{\xi_{1}}\tilde{\vartheta}(\xi)\right]\right\rvert\leq d\int_{{}^{d}}\left\lvert\tilde{\vartheta}(\xi)\right\rvert\left\lvert\xi_{1}\right\rvert\rho(\xi)\mathrm{d}\xi\\ \leq\frac{ld}{2}\int_{{}^{d}}\left\|v-\delta\xi\right\|^{2}\left\lvert\xi_{1}\right\rvert\rho(\xi)\mathrm{d}\xi.$
Since $v=\left\|v\right\|e_{1}$, we have $\left\|v-\delta\xi\right\|^{2}=\left\|v\right\|^{2}+\delta^{2}\left\|\xi\right\|^{2}-2\delta\left\|v\right\|\xi_{1}$. Hence
$\left\lvert\mathbb{E}_{\xi}\left[{}\partial_{\xi_{1}}\tilde{\vartheta}(\xi)\right]\right\rvert\leq\frac{ld}{2}\int_{{}^{d}}\left(\left\|v\right\|^{2}+\delta^{2}\left\|\xi\right\|^{2}\right)\left\lvert\xi_{1}\right\rvert\rho(\xi)\mathrm{d}\xi\\ =\frac{ld}{2}\mathbb{E}_{\xi}\left[{}\left(\left\|v\right\|^{2}+\delta^{2}\left\|\xi\right\|^{2}\right)\left\lvert\xi_{1}\right\rvert\right].$
One can write $\xi=(\xi_{1},\tilde{\xi}^{\top})^{\top}$ with $\xi_{1}\in\mathcal{{}N}\left(0,\frac{1}{d}\right),\tilde{\xi}\in\mathcal{{}N}\left(0,\frac{1}{d}I_{d-1}\right)$. According to Lemma 4.1 and the fact that $\Gamma(z+1)=z\Gamma(z)$,
$\mathbb{E}\left[{}|{\xi_{1}}|\right]=\frac{\sqrt{2}}{\sqrt{d}}\frac{\Gamma\left(1\right)}{\Gamma\left(\frac{1}{2}\right)}=\frac{\sqrt{2}}{\sqrt{d\pi}},\quad\mathbb{E}\left[{}|\xi_{1}|^{3}\right]=\frac{2\sqrt{2}}{d^{\frac{3}{2}}\sqrt{\pi}}\quad\mbox{and}\quad\mathbb{E}\left[{}\left\|\tilde{\xi}\right\|^{2}\right]=\frac{d-1}{d}.$
Therefore,
$\left\lvert\mathbb{E}_{\xi}\left[{}\partial_{\xi_{1}}\tilde{\vartheta}(\xi)\right]\right\rvert$ $\leq\frac{ld}{2}\mathbb{E}_{\xi}\left[{}\left\|v\right\|^{2}\left\lvert\xi_{1}\right\rvert+\delta^{2}\left\lvert\xi_{1}\right\rvert^{3}+\delta^{2}\left\|{\tilde{\xi}}\right\|^{2}\left\lvert\xi_{1}\right\rvert\right]$
$=\frac{ld}{2}\left\|v\right\|^{2}\mathbb{E}_{\xi}\left[{}\left\lvert\xi_{1}\right\rvert\right]+\delta^{2}\frac{ld}{2}\mathbb{E}_{\xi}\left[{}\left\lvert\xi_{1}\right\rvert^{3}\right]+\delta^{2}\frac{ld}{2}\mathbb{E}_{\xi}\left[{}\left\|{\tilde{\xi}}\right\|^{2}\left\lvert\xi_{1}\right\rvert\right]$
$=\frac{ld}{2}\left\|v\right\|^{2}\frac{\sqrt{2}}{\sqrt{d\pi}}+\delta^{2}\frac{ld}{2}\frac{2\sqrt{2}}{d^{\frac{3}{2}}\sqrt{\pi}}+\delta^{2}\frac{ld}{2}\frac{d-1}{d}\frac{\sqrt{2}}{\sqrt{d\pi}}$
$={l}\left\|v\right\|^{2}\frac{\sqrt{d}}{\sqrt{2\pi}}+\delta^{2}{l}\frac{2}{\sqrt{2d\pi}}+\delta^{2}{l}\frac{d-1}{\sqrt{2d\pi}}$
$=l\frac{\sqrt{d}}{\sqrt{2\pi}}\left(\left\|v\right\|^{2}+\delta^{2}\left(1+\frac{1}{d}\right)\right).$
Consequently, one has
$\left\lvert\mathbb{E}_{\xi}\left[{}\partial_{v_{1}}\vartheta(v-\delta\xi)\right]\right\rvert\cdot\left\|v\right\|\leq\frac{1}{\delta}\left\lvert\mathbb{E}_{\xi}\left[{}\partial_{\xi_{1}}\tilde{\vartheta}(\xi)\right]\right\rvert\cdot\left\|v\right\|\leq l\frac{\sqrt{d}\left\|v\right\|}{\delta\sqrt{2\pi}}\left(\left\|v\right\|^{2}+\delta^{2}\left(1+\frac{1}{d}\right)\right),$
so that (4.2) holds.
Recall that $\tilde{\vartheta}(\xi)=\vartheta(v-\delta\xi)$ and $\partial_{\xi_{j}}\tilde{\vartheta}(\xi)=-\delta\partial_{v_{j}}\vartheta(v-\delta\xi),\forall j=1,\ldots,d$. Thus:
$\langle\nabla\vartheta(v-\delta\xi),\delta\xi\rangle=\sum_{j=1}^{d}\partial_{v_{j}}\vartheta(v-\delta\xi)\delta\xi_{j}=-\sum_{j=1}^{d}\partial_{\xi_{j}}\tilde{\vartheta}(\xi)\xi_{j}.$
Note that $\rho(\xi)=\frac{d^{\frac{d}{2}}}{(2\pi)^{\frac{d}{2}}}\exp\left(-\frac{d\left\|\xi\right\|^{2}}{2}\right)$ so that $\nabla\rho(\xi)=-d\rho(\xi)\xi$. Consequently, by partial integration,
$\mathbb{E}_{\xi}\left[{}\langle\nabla\vartheta(v-\delta\xi),\delta\xi\rangle\right]=-\sum_{j=1}^{d}\int_{{}^{d}}\partial_{\xi_{j}}\tilde{\vartheta}(\xi)\xi_{j}\rho(\xi)\mathrm{d}\xi=\sum_{j=1}^{d}\int_{{}^{d}}\tilde{\vartheta}(\xi)\partial_{\xi_{j}}(\xi_{j}\rho(\xi))\mathrm{d}\xi=\int_{{}^{d}}\tilde{\vartheta}(\xi)\text{div}(\rho(\xi)\xi)\mathrm{d}\xi.$
Since $\nabla\rho(\xi)=-d\rho(\xi)\xi$, $\text{div}(\rho(\xi)\xi)=\langle\nabla\rho(\xi),\xi\rangle+d\rho(\xi)=d\rho(\xi)\left(1-\left\|\xi\right\|^{2}\right)$. Therefore,
$\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\vartheta(v-\delta\xi),\delta\xi\rangle\right]\right\rvert$ $\leq\frac{l}{2}\int_{{}^{d}}\left\|v-\delta\xi\right\|^{2}\left\lvert\text{div}(\rho(\xi)\xi)\right\rvert\mathrm{d}\xi=\frac{dl}{2}\mathbb{E}\left[{}\left\|v-\delta\xi\right\|^{2}\left\lvert 1-\left\|\xi\right\|^{2}\right\rvert\right]$
$=\frac{dl}{2}\mathbb{E}\left[{}\left(\left\|v\right\|^{2}+\delta^{2}\left\|\xi\right\|^{2}\right)\left\lvert 1-\left\|\xi\right\|^{2}\right\rvert\right].$
According to (4.1) of Lemma 4.1, we have
$\mathbb{E}\left[{}\left\lvert 1-\left\|\xi\right\|^{2}\right\rvert\right]\leq\sqrt{\mathbb{E}\left[{}\left(1-\left\|\xi\right\|^{2}\right)^{2}\right]}=\sqrt{\mathbb{E}\left[{}1-2\left\|\xi\right\|^{2}+\left\|\xi\right\|^{4}\right]}=\sqrt{\frac{2}{d}},$
and
$\mathbb{E}\left[{}\left\lvert 1-\left\|\xi\right\|^{2}\right\rvert\left\|\xi\right\|^{2}\right]\leq\sqrt{\mathbb{E}\left[{}\left(1-\left\|\xi\right\|^{2}\right)^{2}\right]}\sqrt{\mathbb{E}\left[{}\left\|\xi\right\|^{4}\right]}=\sqrt{\frac{2}{d}}\sqrt{1+\frac{2}{d}}\leq\sqrt{\frac{2}{d}}\left(1+\frac{1}{d}\right).$
Therefore,
$\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\vartheta(v-\delta\xi),\delta\xi\rangle\right]\right\rvert\leq l\sqrt{\frac{d}{2}}\left\|v\right\|^{2}+l\delta^{2}\sqrt{\frac{d}{2}}\left(1+\frac{1}{d}\right)\\ =l\sqrt{\frac{d}{2}}\left(\left\|v\right\|^{2}+\delta^{2}\left(1+\frac{1}{d}\right)\right),$
which implies (4.3). Summing (4.2) and (4.3) up, one can readily get
$\left\lvert\mathbb{E}_{\xi}\left[{}\langle\nabla\varepsilon(x-\delta\xi),x-\delta\xi-x^{*}\rangle\right]\right\rvert\leq l\sqrt{\frac{d}{2}}\left(\frac{\left\|x-x^{*}\right\|}{\delta\sqrt{\pi}}+1\right)\left(\left\|x-x^{*}\right\|^{2}+\delta^{2}\left(1+\frac{1}{d}\right)\right),$
which implies (4.4), and this completes the proof. ∎
A.2. Proof of Lemma 4.3
Proof.
Define the stopping time
$\tau:=\inf\{t\mid X_{t}<\ell\}.$
[…]
By taking expectation with respect to $X_{0},\ldots,X_{t}$ on both sides one can get
(A.1) $\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t\}\right]\leq\theta\mathbb{E}\left[{}X_{t}\mathds{1}\{\tau>t\}\right].$
Note that $\mathds{1}\{\tau>t\}=\mathds{1}\{\tau>t+1\}+\mathds{1}\{\tau=t+1\}$, so that
$\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t\}\right]$ $=\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t+1\}\right]+\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau=t+1\}\right]$
(A.2) $\geq\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t+1\}\right]-b\mathbb{P}\left[{}\tau=t+1\right].$
Combining (A.1) and (A.2) together obtains
$\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t+1\}\right]\leq\theta\mathbb{E}\left[{}X_{t}\mathds{1}\{\tau>t\}\right]+b\mathbb{P}\left[{}\tau=t+1\right].$
Then we have
(A.3) $\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t+1\}\right]\leq\theta^{t+1}\mathbb{E}\left[{}X_{0}\mathds{1}\{\tau>0\}\right]+b\sum_{k=1}^{t+1}\theta^{t+1-k}\mathbb{P}\left[{}\tau=k\right].$
Note that
(A.4) $\sum_{k=1}^{t+1}\theta^{t+1-k}\mathbb{P}\left[{}\tau=k\right]$ $=\sum_{k=1}^{t+1}\theta^{t+1-k}\left(\mathbb{P}\left[{}\tau>k-1\right]-\mathbb{P}\left[{}\tau>k\right]\right)$
$=\sum_{k=0}^{t}\theta^{t-k}\mathbb{P}\left[{}\tau>k\right]-\sum_{k=1}^{t+1}\theta^{t+1-k}\mathbb{P}\left[{}\tau>k\right]$
$=(1-\theta)\sum_{k=0}^{t}\theta^{t-k}\mathbb{P}\left[{}\tau>k\right]+\theta^{t+1}\mathbb{P}\left[{}\tau>0\right]-\mathbb{P}\left[{}\tau>t+1\right].$
Moreover, according to (4.5) one has $B=\mathbb{E}\left[{}X_{0}\mathds{1}\{X_{0}\geq\ell\}\right]=\mathbb{E}\left[{}X_{0}\mathds{1}\{\tau>0\}\right]\geq\mathbb{E}\left[{}\ell\mathds{1}\{\tau>0\}\right]=\ell\mathbb{P}\left[{}\tau>0\right]$, so that
(A.5) $\mathbb{P}\left[{}\tau>0\right]\leq\dfrac{B}{\ell}.$
Therefore, by substituting (A.4) and (A.5) into (A.3), we have
(A.6) $\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t+1\}\right]$ $\leq\theta^{t+1}B+b\left((1-\theta)\sum_{k=0}^{t}\theta^{t-k}\mathbb{P}\left[{}\tau>k\right]+\theta^{t+1}\dfrac{B}{\ell}-\mathbb{P}\left[{}\tau>t+1\right]\right)$
$\leq\theta^{t+1}\left(1+\frac{b}{\ell}\right)B-b\mathbb{P}\left[{}\tau>t+1\right]+(1-\theta)b\sum_{k=0}^{t}\theta^{t-k}\mathbb{P}\left[{}\tau>k\right].$
On the other hand, it holds that $\mathbb{E}\left[{}X_{t+1}\mathds{1}\{\tau>t+1\}\right]\geq\ell\mathbb{P}\left[{}\tau>t+1\right]$, which together with (A.6) implies
(A.7) $\mathbb{P}\left[{}\tau>t+1\right]\leq\theta^{t+1}\frac{B}{\ell}+\frac{(1-\theta)b}{b+\ell}\sum_{k=0}^{t}\theta^{t-k}\mathbb{P}\left[{}\tau>k\right].$
In the following, we prove by conduction that
(A.8) $\mathbb{P}\left[{}\tau>t\right]\leq\left(\frac{b+\theta\ell}{b+\ell}\right)^{t}\frac{B}{\ell},\quad\forall 0\leq t\leq M.$
Note that (A.8) trivially holds for $t=0$ according to (A.5). Now, assume that (A.8) is verified for all $t\leq n$ with $n\geq 0$. Then by using (A.7) one can get
(A.9) $\mathbb{P}\left[{}\tau>n+1\right]\leq\theta^{n+1}\frac{B}{\ell}+\frac{(1-\theta)b}{b+\ell}\sum_{t=0}^{n}\theta^{n-t}\left(\frac{b+\theta\ell}{b+\ell}\right)^{t}\frac{B}{\ell}.$
Note that
$\sum_{t=0}^{n}\left(\frac{b+\theta\ell}{\theta b+\theta\ell}\right)^{t}=\frac{\left(\frac{b+\theta\ell}{\theta b+\theta\ell}\right)^{n+1}-1}{\left(\frac{b+\theta\ell}{\theta b+\theta\ell}\right)-1}=\frac{\theta b+\theta\ell}{(1-\theta)b}\left(\left(\frac{b+\theta\ell}{\theta b+\theta\ell}\right)^{n+1}-1\right).$
Thus we can get from (A.9) that
$\mathbb{P}\left[{}\tau>n+1\right]$ $\leq\theta^{n+1}\frac{B}{\ell}+\frac{(1-\theta)b}{b+\ell}\theta^{t}\sum_{t=0}^{n}\left(\frac{b+\theta\ell}{\theta b+\theta\ell}\right)^{t}\frac{B}{\ell}$
$=\theta^{t+1}\frac{B}{\ell}+\theta^{t+1}\left(\left(\frac{b+\theta\ell}{\theta b+\theta\ell}\right)^{n+1}-1\right)\frac{B}{\ell}=\left(\frac{b+\theta\ell}{b+\ell}\right)^{n+1}\frac{B}{\ell}.$
Consequently, (A.8) holds. Therefore,
$\mathbb{P}\left[{}\tau\leq M\right]\geq 1-\left(\frac{b+\theta\ell}{b+\ell}\right)^{M}\frac{B}{\ell}.$
Since the event $\tau\leq M$ is equivalent to the existence of $t\leq M$ such that $X_{t}<\ell$, the Lemma is proved. ∎